### v1

# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

Abstract
With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that evaluates both spatial and textual relevance, have found many real-life applications. Existing geo-textual indexes for TkQs use traditional retrieval models like BM25 to compute text relevance and usually exploit a simple linear function to compute spatial relevance, but its effectiveness is limited. To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues. To the best of our knowledge, there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST). LIST is featured with two novel components. Firstly, we propose a lightweight and effective relevance model that is capable of learning both textual and spatial relevance. Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. Two key challenges in building an effective and efficient index are the absence of high-quality labels and unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
I INTRODUCTION
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps [1]; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to retrieve geo-textual objects with texts relevant to the keywords ‘delicious Italian restaurants’ and located near the query’s location, and the result is a list of $k$ geo-textual objects ranked by a relevance model considering both spatial and textual relevance.
Spatial keyword queries have applications in many scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions. To meet diverse user needs, various spatial keyword queries have been introduced [2, 3, 4]. Among them, the Top-k KNN Spatial Keyword Query (TkQ[2] retrieves top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. According to the experimental evaluation [5], TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query [4], which retrieves the geo-textual objects containing all the query keywords and rank them according to their spatial relevance (or proximity) to the query location. Most of the existing studies on spatial keyword queries focus on the improvement in the efficiency of handling spatial keyword queries. As such, many indexes [6] and corresponding query processing algorithms have been developed.
Motivations. The existing indexes [2, 7, 8, 9] designed for the TkQ query typically compute textual relevance using traditional retrieval models such as BM25 [10] and TF-IDF [11], which rely on exact word matching to evaluate textual relevance. These traditional models often suffer from the word mismatch issue, i.e., synonyms or paraphrases that consist of completely different tokens may convey the same meanings, which limits their effectiveness (as to be shown in Section III-B). For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models, although they are relevant to the query.
As for spatial relevance, existing indexes [2, 7, 8, 9] for the TkQ query typically compute spatial relevance by a linear function of the distance between the query location and the location of a geo-textual object. The implicit assumption behind this practice is that the spatial preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we find it does not hold on our real-life datasets. From our dataset, we find that the spatial preference for geo-textual objects is not a linear function of distance, but exhibits a significant (non-linear) decrease with the increase of distance (detailed in Section III-B). This motivates us to design a more effective spatial relevance module. A straightforward approach would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (shown in Section V-E).
To improve effectiveness, several deep learning based relevance models [5, 12, 13] propose to to compute textual relevance based on word embeddings. For example, DrW [5] employs embeddings generated from the Bert model, identifying top-k relevant terms from geo-textual object description for each query keyword based on the embeddings, and uses an attention mechanism to aggregate scores of each query keyword. Although these approaches could achieve better effectiveness, they have very high query latency for computing the relevance score since they rely on embeddings and neural networks (detailed in Section III-B). To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models.
Existing geo-textual indexes are developed based on traditional retrieval models such as TF-IDF. For instance, IR-Tree [2] augments each R-Tree [14] node with a pseudo-inverted file including the words of all objects inside that node. These indexes cannot be used to handle embedding based spatial keyword queries, which represent the semantic meaning of query keywords with a high-dimensional embedding. Although Approximate Nearest Neighbor Search (ANNS) indexes [15, 16, 17] are designed for embedding retrieval, none of these ANNS indexes take into account the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section V-B). It is an open problem to incorporate spatial factors into ANNS indexes to support embedding based spatial keyword queries.
Objectives. To address the aforementioned issues, this work aims to (1) design an embedding based lightweight yet effective spatio-textual relevance model as the ranking function, and (2) design a novel index to speed up high-dimensional embedding based spatial keyword queries.
To achieve the two objectives, we propose a novel solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries. We denote this retriever by LIST, which is featured with two new components. First, LIST develops a lightweight and effective relevance model as the ranking function, which comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module.
Second, LIST introduces a novel machine learning based ANNS index. For both embedding based spatial keyword queries and geo-textual objects, our proposed index utilizes the learning-to-cluster technique, which was developed for image clustering in the field of computer vision, to cluster relevant geo-textual objects and queries together and build an index. Existing learning-to-cluster techniques  [18, 19] rely on partial pairwise similar/dis-similar relationships of images for training, where the pairwise labels are easy to obtain for images. However, such pairwise positive labels for queries and geo-textual objects in our problem are very sparse and high-quality negative labels are absent, which are important for learning-to-cluster. Additionally, when the number of clusters exceeds the number of image classes, the existing learning-to-cluster technique will produce a highly skewed cluster distribution [18], which will hurt the efficiency. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. These informative negative labels combined with ground-truth positive (relevant) labels (e.g., click-through or human annotation, detailed in Section V-A) enable our index to learn both spatial and textual information, thereby clustering relevant objects and queries together while separating irrelevant objects and queries.
The contributions of this work are summarized as follows:
  • New Index. We develop a novel machine learning based ANNS index that can speed up embedding based spatial keyword queries. Different from existing ANNS indexes, it extends the learning-to-cluster technique to cluster relevant objects and queries together. To build an effective and efficient index, we propose a novel pseudo-label generation approach. To the best of our knowledge, this is the first index tailored for embedding based spatial keyword queries. Additionally, this is the first geo-textual index to utilize a neural network for retrieval without relying on an explicit tree structure and associated heuristic algorithms.
  • New Relevance Model. We develop a lightweight and effective relevance model. It is featured with a novel learning-based spatial relevance module to learn spatial relevance adaptively, a new weight learning mechanism to determine the weights between textual and spatial relevance, and a dual-encoder module to compute textual relevance.
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets. Experimental results show that LIST outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline. Our implementation is released at https://github.com/Heisenberg-Yin/LIST.
II RELATED WORK
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries [2, 3, 4] are proposed. Among them, Top-k KNN Spatial-Keyword Query (TkQ[2] aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial relevance and text relevance. The earlier work [2] computes the textual relevance with traditional retrieval models like BM25 [10] and TF-IDF [11] and exploits a simple linear function for spatial relevance. However, these traditional methods have limited effectiveness.
To enhance the effectiveness, several deep learning based methods [13, 5, 12] have been developed for query-POI matching, which is essentially spatial keyword query. PALM [13] considers geographic information by using location embedding techniques and combines it with semantic representations for query-POI matching. DrW [5] computes the deep textual relevance on the term level of query and geo-textual objects and uses the attention mechanism to aggregate scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual relevance and spatial relevance. MGeo [12] employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality. MGeo aligns these two modalities into the same latent space and computes relevance scores based on the produced representations. However, these deep learning based methods focus on improving the effectiveness, but ignore the efficiency issue.
Spatio-Textual Indexes. Although various spatio-textual indexes [3, 4, 7, 2, 8, 9, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] have been designed to efficiently answer spatial keyword queries, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for high-dimensional embedding based retrieval models used by deep learning based methods.
[…]
Different from previous spatio-textual indexes developed to accelerate the search for top-k objects with traditional relevance models (which is called sparse retrieval), we aim to design an index to accelerate the search for top-k objects with embedding based relevance models (which is called dense retrieval), which is still an open problem for geo-textual data.
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models [34]. The first category of models (e.g., ARC-II [35] and MatchPyramid [36]) calculates the word-level similarity between queries and documents for the final relevance scores. DrW [5] belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM [37]) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of the dual-encoder architecture [38, 39], which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval [40, 41, 42, 43, 44]. For the second category of methods, to support efficient retrieval, the learned embeddings are usually indexed for Approximate Nearest Neighbor Search (ANNS) [34].
ANNS indexes are developed to expedite the top-k search on high-dimensional data. These techniques can be broadly categorized into two types. The first type focuses on searching a subset of database. The representative methods include inverted file index (IVF) based methods [45, 46, 47], hashing-based methods [48, 49, 50, 51], and graph-based methods [52, 17]. These ANNS indexes rely on heuristic algorithms or centroids to find candidates. Specifically, Inverted File Index (IVF) [46] partitions data into clusters using the k-means algorithm and routes queries to clusters based on the query’s distance to the clusters’ centroids. Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index [52] construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings. Hashing-based methods like Locality-Sensitive Hashing (LSH) index [53] generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates. The other type aims to accelerate the search process itself, such as quantization-based methods [46, 54, 55, 56, 54].
Our method belongs to the first category and is orthogonal to the second category. Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatio-textual relevance models. In contrast, existing ANNS indexes are only for textual relevance. (2) Instead of using heuristic algorithms to generate candidates, our proposed index learns from query-object pairs in a data-driven manner and employs neural networks for data partitioning and query routing. Notably, our proposed index utilizes a novel learning-to-cluster technique to group geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
III PROBLEM STATEMENT AND MOTIVATIONS
III-A Preliminary
We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual document description $o.doc$. The location description $o.loc$ is a two-dimensional GPS coordinate composed of latitude and longitude. The document description $o.doc$ reflects the textual content. The Top-k KNN Spatial-Keyword Query (TkQ[2] is defined as follows.
Top-k KNN Spatial-Keyword Query (TkQ): Given a query $q=\langle loc,doc,k\rangle$, where $q.doc$ denotes the query keywords, $q.loc$ is the query location, and $q.k$ is the number of returned objects, we aim to retrieve $k$ objects with the highest relevance scores $ST(q,o)$:
$ST(q,o)=(1-\alpha)\times SRel(q.loc,o.loc)+$ (1)
$\alpha\times TRel(q.doc,o.doc).$
[…]
III-B Data Analysis and Motivations
(a) Word Mismatch Analysis
(b) Spatial Relevance Analysis
Figure 1: Figure 1(a) shows the percentage distribution of ground-truth positive query-object pairs based on the number of matching terms on the Beijing dataset. Figure 1(b) compares the CDF of spatial distance for ground-truth positive query-object pairs and the linear distribution on the Beijing dataset.
Word Mismatch. As discussed earlier, existing indexes [2, 7, 9] utilize traditional ranking methods, such as BM25 [10] and TF-IDF [11], to compute textual relevance. These models often suffer from the word mismatch issue, which reduces their effectiveness. As shown in Figure 1(a), in the Beijing dataset, nearly 20% of the ground-truth query-object positive (relevant) pairs have no overlap of words. Since these traditional models rely on exact word matching to compute text relevance, relevant objects may receive a low or even zero textual relevance from these models. For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero.
Spatial Relevance. Previous studies [2, 5] typically compute spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption for the linear formulation is that the user preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not really hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario.
Efficiency. To address this issue, several recent studies [5, 13, 57] have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness. However, these methods rely on word embeddings and deep neural networks to compute textual relevance, resulting in high querying latency. For instance, DrW [5] needs to identify the top-k relevant words from each object for each query word, and then aggregate these word pairs using an attention mechanism to compute the textual relevance. On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in [5]. This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as rerankers for a small number of retrieved objects.
Incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. On the one hand, the index techniques developed for spatial keyword queries are designed based on the traditional retrieval models such as BM25 and cannot be used to handle embedding-based spatial keyword queries. On the other hand, existing ANNS indexes, which are discussed in Section II, do not consider spatial relevance. It is an open problem of incorporating spatial relevance into the ANNS indexes to support spatial keyword queries. Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section V-B).
III-C Problem Statement
To address these issues, we aim to (1) design a lightweight and effective embedding-based spatio-textual relevance model as the ranking function, and (2) develop an index that can accelerate the top-k search process of embedding based spatial keyword queries.
IV PROPOSED RETRIEVER (LIST)
Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. LIST consists of a relevance model (shown in yellow) and an index (shown in green).
IV-A Overview
As discussed in the introduction, LIST is featured with a novel machine learning based ANNS index and a new relevance model. Notably, our index is applicable to any textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase. During the training phase, we first train our relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to a cluster or a subset of clusters that have the highest probabilities. In the cluster, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
IV-B The Proposed Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s great success in document retrieval [38], we employ a dual-encoder module to encode both the query keywords and the document into two separate embeddings and calculate textual relevance using the two embeddings, which is used in spatial keyword query for the first time. Compared with previous relevance models [5] which rely on word embeddings and complex nerual networks, this approach is more efficient.
This module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert [58]. The encoder takes the textual content of a query or object as input, captures interactions between terms by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional dense vector. This process is formulated as:
$o.emb=E_{o}(o.doc;\theta_{o}),\>o.emb\in\mathbb{R}^{d},$ (2)
$q.emb=E_{q}(q.doc;\theta_{q}),\>q.emb\in\mathbb{R}^{d}.$
where $E_{o}(\textperiodcentered;\theta_{o})$ denotes the object encoder parameterized with $\theta_{o}$ and $E_{q}(\textperiodcentered;\theta_{q})$ denotes the query encoder parameterized with $\theta_{q}$. Then the text relevance score can be calculated by the inner product between the $q.emb$ and $o.emb$,
$TRel(q.doc,o.doc)=q.emb\cdot o.emb.$ (3)
Spatial Relevance Learning. To fit the real-world scenario described in Section III-B, it is essential to assign higher spatial relevance to nearby objects and lower spatial relevance to distant objects. Therefore, we propose a novel learning-based spatial relevance module to better capture spatial relevance, where $S_{in}=1-SDist(q.loc,o.loc)$ is taken as input.
The user’s preference for geo-textual objects increases as the distance decreases and exhibits a stepwise decline as the distance increases, i.e., users tend to visit nearby objects and become less tolerant as the distance increases. To illustrate this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, a new spatial relevance module is designed, which is a learnable monotonically increasing step function, to align with this real-world characteristic.
Our learning-based spatial relevance module consists of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$. Here, $\mathcal{T}$ is used to ensure that the learned function is a step function and $w_{s}$ is employed to learn and estimate the spatial relevance. Specifcially, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance. $\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value. For example, when $t=100$, $\mathcal{T}$ is $[0.0,0.01,\cdots,0.99,1.0]$. The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values. Note that weights in $w_{s}$ are learned from the training data. When the input $S_{in}$ exceeds the value of $\mathcal{T}[i]$, then the spatial relevance increases by $act(w_{s}[i])$, where $act$ is an activation function to ensure $act(w_{s}[i])$ remain non-negative. This process ensures that the output $SRel$ exhibits a step increase as the input $S_{in}$ increases. The learned spatial relevance is computed as below:
$SRel(q.loc,o.loc)=act(w_{s})\cdot\mathbb{I}(S_{in}\geq\mathcal{T}[i]),$ (4)
where $SRel(q.loc,o.loc)$ is the learned spatial relevance. $\mathbb{I}\in\{0,1\}^{t\times 1}$ is an indicator array. $\mathbb{I}[i]=1$ if $S_{in}\geq$ $\mathcal{T}[i]$; otherwise 0. The sum of the step increment is conducted by an inner product between the indicator array $\mathbb{I}\in\{0,1\}^{t\times 1}$ and the learnable weight after activation $act(w_{s})\in\mathbb{R}^{1\times t}$.
During the query phase, we extract the weights in $w_{s}$ from the module and store them as an array $\hat{w_{s}}$ for faster inference. $\hat{w_{s}}$ is constructed as $\hat{w_{s}}[i]=\sum\limits_{0}^{i}act(w_{s}[i])$. When computing the spatial relevance, we get the input $S_{in}$. Since the threshold value grows uniformly by $\frac{1}{t}$, we can determine the number of threshold values exceeded by the input as $\lfloor S_{in}*t\rfloor$, where $\lfloor$.$\rfloor$ . ⌋ indicates a floor function and is utilized to truncate a real number to an integer. This also corresponds to the sum of the values of weights, which is the spatial relevance score. This process is formulated as:
$SRel(q.loc,o.loc)=\hat{w_{s}}[\lfloor S_{in}*t\rfloor].$ (5)
Hence, during the query phase, the time complexity of computing spatial relevance is O(1), which is efficient.
Adaptive Weight Learning. The latest study [5] has showed the importance of weight learning in improving ranking effectiveness. We rely on the dual-encoder architecture to obtain embedding of query $q.emb$ to compute textual relevance. In addition, our empirical experiments show that the location embedding technique would introduce additional computational costs. Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine an optimal weight based on $q.emb$, formulated as follows:
$w_{st}=MLP(q.emb),\>w_{st}\in\mathbb{R}^{1\times 2}.$ (6)
Similar to the Equation 1, the final relevance score between the query and the object is calculated as:
$ST(q,o)=w_{st}\cdot[TRel(q.doc,o.doc),SRel(q.loc,o.loc)]^{T}.$ (7)
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section V-A). For the negative (irrelevant) geo-textual objects, we choose a subset of hard negative objects for training. Specifically, following previous work [5], for each training query, in each training epoch, we randomly pick $b$ hard negative samples from this set. We optimize the loss function as the negative log-likelihood of the positive object:
$\begin{split}L_{\text{model}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,b}^{-})\\ &=-log\frac{e^{ST(q_{i},o_{i}^{+})}}{e^{ST(q_{i},o_{i}^{+})}+\sum_{j=1}^{b}e^{ST(q_{i},o_{i,j}^{-})}}.\end{split}$ (8)
In addition, we also utilize the in-batch negatives strategy [59] to further enhance training efficiency.
IV-C The Proposed Index
Figure 3 illustrates our proposed index, which takes geo-location and textual embeddings as input. It uses a Multi-Layer Perception (MLP) to produce the $c$-cluster probability distribution and employs training strategies to ensure relevant queries and objects have similar probability distributions.
Feature Construction. We utilize geo-location and textual embedding to construct a consistent input representation for both objects and queries, which is illustrated in Figure 3. The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding, which is converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module. The geo-location $loc=\langle lat,lon\rangle$ are transformed into the following features $\langle\hat{lat},\hat{lon}\rangle$ as below:
$\hat{lat}=\frac{lat-lat_{min}}{lat_{max}-lat_{min}},\hat{lon}=\frac{lon-lon_{min}}{lon_{max}-lon_{min}},$ (9)
where $lat_{min}$ ($lon_{min}$) represents the lowest latitude (longitude) in dataset $D$, $lat_{max}$ ($lon_{min}$) denotes the highest latitude (longitude) in dataset $D$. Then the input representation is formulated as
$x=[emb,\hat{lat},\hat{lon}].$ (10)
Here $x$ is the input feature vector for the cluster classifier. Although several region embedding techniques exist [60], we find that this simple approach is sufficiently effective. Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
$Prob=Softmax(MLP(x)),\>Prob\in\left[0,1\right]^{c\times 1},$ (11)
where $c$ is a hyperparameter that indicates the desired number of clusters we aim to obtain, and $Prob$ represents the predicted $c$-cluster probability distribution. The MLP is shared between queries and objects. The $c$-cluster probability distribution of object $o$ is denoted as $Prob_{o}$ and that of query $q$ is represented as $Prob_{q}$. The cluster classifier actually serves as a relevance classifier, grouping identified relevant queries and objects into the same cluster. For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ forms a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
Figure 3: The illustration of the index.
Pseudo-Label Generation. To train the cluster classifier, both positive and negative pairwise labels are required. We utilize the ground-truth query-object relevant label as the positive pairwise label for training. However, we lack high-quality negative pairwise labels. Randomly selecting negative objects will lead to overfitting and all objects are grouped together (as detailed later). Therefore, we propose a novel pseudo pairwise negative label generation method.
The relationship between query $q$ and object $o$ is denoted as $s(q,o)$. If $q$ and $o$ are relevant , $s(q,o)=1$; otherwise, $s(q,o)=0$, which are then used as labels for training. We leverage ground-truth positive labels as pairwise positive labels in the training, as shown below:
$\text{pos}_{q}=\{o;o\in D,s(q,o)=1\},$ (12)
where the positive object set of query $q$ is denoted as $\text{pos}_{q}$.
The negative object set $\text{neg}_{q}$ of query $q$ is generated by the relevance model. Given a query $q$, we employ the relevance model to calculate the relevance score for all objects. Then we adaptively select $\text{neg}_{q}$ according to two hyperparameters $neg_{start}$ and $neg_{end}$, as shown below:
$\text{neg}_{q}=\text{argsort}_{o\in D}ST(q,o)[neg_{start}:neg_{end}],\;s(q,o)=0,$ (13)
where $ST(q,o)$ is the relevance score between query $q$ and object $o$ produced by the relevance model. Note that positive query-object pairs are excluded as indicated by the filter condition $s(q,o)=0$.
This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy [44, 61]. What sets our approach apart from existing studies is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples. This adjustment strikes the trade-off between the effectiveness and efficiency.
Decreasing $neg_{start}$ leads to a set of harder negative objects being chosen to train the model. Consequently, the classifier is more effective in distinguishing between positive objects and hard negative objects, which facilitates the clustering of relevant queries and objects, while effectively segregating the irrelevant ones. Empirically, under this setting, only the very relevant queries and objects are grouped into the same cluster. The reduced number of objects within a query’s cluster leads to higher efficiency. In contrast, when using a large $neg_{start}$, the index can not learn useful information, leading to a scenario where all objects tend to be clustered together.
On the other hand, when our index is trained on a set of harder negative objects, it might also exclude some positive ones and reduce its effectiveness. Thus, the choice of $neg_{start}$ strikes a trade-off between effectiveness and efficiency. This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques [18, 19].
Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function [62] to train the MLP, as described below:
$\begin{split}L_{\text{Index}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,m}^{-})\\ &=log(\hat{s}(q_{i},o_{i}^{+}))+\sum\limits_{j=1}^{m}log(1-\hat{s}(q_{i},o_{i,j}^{-})),\end{split}$ (14)
where $\hat{s}(q_{i},o_{j})=Prob^{T}_{q_{i}}\cdot Prob_{o_{j}}$, $o_{i}^{+}\in\text{pos}_{q}$ and $o_{i,j}^{-}\in\text{neg}_{q}$. Typically, we randomly select one positive object $o_{i}^{+}$ from the positive object set $\text{pos}_{q}$ and $m$ negative objects from the negative object set $\text{neg}_{q}$ in each training epoch.
As we aim at grouping the relevant pairs of query and object into the same cluster while the irrelevant pairs are into distinct clusters, the queries and objects are jointly incorporated. This is because the MLP is shared between objects and queries. In this manner, the complex relationships between queries and objects can be well learned in the process of index construction. Figure 3 also demonstrates this process, where relevant pairs share similar output and irrelevant ones have different output.
Learning to Partition and Route. During the indexing phase, each object $o$ is partitioned to the cluster with the highest probability according to $Prob_{o}$. Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which can act as an inverted file for these objects. Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which are utilized to calculate the relevance score for incoming queries. During the query phase, a given query $q$ is directed to the cluster that has the highest probability. Subsequently, the relevance scores between $q$ and all objects in the cluster are calculated and the top-k objects are selected as result objects.
An alternative way is routing queries (objects) to $cr$ clusters with the highest probabilities based on $Prob_{q}$ ($Prob_{o}$). Although this might boost query effectiveness by considering more objects in different clusters, it sacrifices efficiency as more objects need to be calculated for a query, leading to an accuracy-efficiency trade-off.
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. Note that most existing ANNS indexes are static [48, 16, 17]. When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. Our index follows this approach. Apparently, the time costs of insertion and deletion operations are negligible.
Cluster Evaluation. After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$. We use validation queries to evaluate the quality of the clusters. The validation queries are fed into the trained index and routed to a cluster. $C_{i}^{q}$ denotes a list containing the validation queries routed to $C_{i}$, $C_{i}^{o}$ denotes a list containing objects partitioned to $C_{i}$, and ||||| | indicates the size of a list. We proceed to introduce two metrics to evaluate the quality of the clusters.
The first metric evaluates the precision of $C_{i}$, denoted as $P(C_{i})$, representing the degree to which queries are aligned with their corresponding positive (relevant) objects in the same cluster, which is defined below:
$P(C_{i})=\frac{1}{|C_{i}^{q}|}\sum\limits_{q_{j}\in C_{i}^{q}}\frac{|\text{pos}_{q_{j}}\cap C_{i}^{o}|}{|\text{pos}_{q_{j}}|}.$ (15)
Building upon this, we compute the average precision across all clusters, denoted as $P(C)$, which is defined as below:
$P(C)=\frac{1}{\sum\limits_{C_{i}\in C}|C_{i}^{q}|}\sum_{C_{i}\in C}P(C_{i})*|C_{i}^{q}|.$ (16)
Intuitively, a higher $P(C)$ indicates that the index is more effective.
In addition, we also consider concerns of efficiency. For this purpose, we introduce another metric, the Imbalance Factor (IF) [63], which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$. The $IF(C)$ is formulated as:
$\text{IF}(C)=\frac{\sum|C_{i}|^{2}}{(\sum|C_{i}|)^{2}},$ (17)
where $\text{IF}(C)$ is minimized when $|C_{1}|=|C_{2}|=\cdots=|C_{c}|$ according to the Cauchy-Schwarz Inequality. A higher imbalance factor indicates a more uneven distribution of clusters. When most objects are concentrated in a few large clusters, the imbalance factor increases significantly, which is undesirable for our task. Overall, our goal is to achieve high-quality clustering results characterized by higher $P(C)$ and lower $\text{IF}(C)$. In our index, with proper hyperparameters, our training process of the cluster classifier is able to obtain high-quality clusters. In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section V-E).
IV-D Procedures and Analyses of LIST
Input: A geo-textual object dataset $D$, a relevance model $R$, a index $I$, training TkQs set $Q_{\text{train}}$, and incoming TkQs set $Q$
Output: The response $Res_{q}$ for each query $q\in Q$
1 // Training Phase: $Train(Q_{\text{train}},D,R,I)$
2 Train $R$ by $Q_{\text{train}}$, $D$ based on Eqaution 8;
3 Employ $R$ to generate pseudo-labels based on Equation 13;
4 Employ pseudo-labels to train $I$ based on Equation 14;
5 // Indexing Phase: $Indexing(D)$
6 for $o\in D$ do
7       Transform $o.doc$ to $o.emb$ based on Equation 2;
8       Transform $o.emb$, $o.loc$ to $x_{o}$ based on Equation 10;
9       Generate $Prob_{o}$ by $MLP$ based on Equation 11;
10       Parition $o$ to cluster $C_{i}$ based on $Prob_{o}$;
11
12 // Query Phase: $Search(q,C,R,I)$
13 for $q\in Q$ do
14       Transform $q.doc$ to $q.emb$ based on Equation 2;
15       Transform $q.emb$, $q.loc$ to $x_{q}$ based on Equation 10;
16       Generate $Prob_{q}$ by $I$ based on Equation 11;
17       Route $q$ to cluster $C_{i}$ based on $Prob_{q}$;
18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
19       return $Res_{q}$;
20
Algorithm 1 Procedures of LIST.
Procedures of LIST. The detailed procedures of LIST are summarized at Algorithm 1. In the training phase, we train the proposed relevance model and index (lines 2-4). After that, in the indexing phase, we assign all objects to their corresponding clusters as inverted files (lines 6-10). In the query phase, given a new query $q$, we extract its features and then route it to a cluster $C_{i}$ by the index (lines 13-16). Consequently, we calculate the relevance score between $q$ and each object $o$ in that cluster by the relevance model, and the top-$k$ objects with the highest scores are returned to answer query $q$ (lines 17-18).
Complexity Analysis. Now we analyze the time and space complexity of LIST. Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the number of layers in the MLP of the index is $l$. The embeddings $o.emb$ of objects are generated in advance.
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section V-E), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3. $O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
The space complexity of LIST is $O((l-1)d+dc)+n(d+2)$. $O((l-1)d+dc)$ is the size of the MLP used by our index. $O(nd)$ represents the space required for storing pre-computed $o.emb$, and $O(2n)$ denotes the storage cost for geo-locations.
V EXPERIMENTS
TABLE I: Datasets Statistics.
Beijing
Shanghai
Geo-Glue
Number of Pois
122,420
116,859
2,849,754
Number of Queries
168,998
127,183
90,000
Number of Records
233,343
182,634
90,000
Number of Train Queries
136,890
103,019
50,000
Number of Val Queries
15,209
11,446
20,000
Number of Test Queries
16,899
12,718
20,000
Number of Train Records
189,027
148,017
50,000
Number of Val Records
21,034
16,492
20,000
Number of Test Records
23,282
18,125
20,000
We evaluate the effectiveness and efficiency of our proposed solution in answering TkQs by comparing with state-of-the-art methods on three real-world datasets. We aim to answer the following research questions:
  • RQ1: Does our proposed retriever LIST and our proposed relevance model outperform previous state-of-the-art methods in terms of effectiveness?
  • RQ2: How well does the efficiency of our proposed retriever LIST compare with state-of-the-art methods?
  • RQ3: How does LIST perform on datasets and training datasets of varying sizes?
  • RQ4: What are the impacts of our proposed index and different modules in the proposed LIST?
TABLE II: Comparison of relevance models across three datasets by brute-force search.
Beijing Shanghai Geo-Glue
Recall NDCG Recall NDCG Recall NDCG
@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
TkQ
0.5740
0.5283
0.4111
0.3302
0.6746
0.6380
0.5044
0.4009
0.5423
0.5023
0.3847
0.3051
PALM
0.3514
0.3098
0.2077
0.1343
0.4617
0.4023
0.2065
0.1223
N/A
N/A
N/A
N/A
DrW
0.6968
0.6316
0.4814
0.3791
0.7689
0.7159
0.5394
0.4114
N/A
N/A
N/A
N/A
DPR
0.5078
0.4269
0.2892
0.2209
0.5143
0.4161
0.2505
0.1721
0.7626
0.7249
0.6014
0.5079
${}_{D}$
0.7555 0.6845 0.5181 0.4215 0.7986 0.7494 0.5769 0.4568 0.7963 0.7565 0.6211
0.5202
MGeo
N/A
N/A
N/A
N/A
N/A
N/A
N/A
N/A
0.7049
N/A
N/A
0.5270
LIST-R
0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
7.95%
10.22%
14.12%
18.36%
4.69%
5.73%
11.71%
18.14%
5.73%
6.18%
6.26%
11.70%
V-A Experimental Setup
Datasets. We utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. The Beijing and Shanghai datasets [5] are provided by Meituan, a Chinese retail services platform. Users submit queries consisting of location and keywords. The clicked Points of Interest (POIs) recorded in the query log are regarded as the ground truth answers to these queries.Note that the click-through data recorded in the search log is widely used as ground-truth for information retrieval evaluation as explicit feedback such as ratings is challenging to collect and implicit feedback such as click-through data can be tracked automatically and is thus much easier to collect for content providers [64]. Both datasets are anonymized to protect user privacy and more details can be found here [5]. The Geo-Glue dataset [65, 12] is released by DAMO Academic. The POIs in the dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
Note that, due to privacy considerations, the release of query log datasets from the industry is restricted. As a result, to our knowledge, no other datasets containing ground-truth query results have been released, except for the three datasets used in our experiment. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section V-D)
Dataset Split. The statistics of datasets are stated in Table I, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects. For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy [5], where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set. From the training set, we randomly choose 10% data as the validation set to tune hyperparameters. For the Geo-Glue dataset [12], we follow the provided splits for training, validation, and test data.
Effectiveness metric. Following previous studies [12, 5], we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness. Recall@$k$ evaluates the proportion of positive objects contained in the top-$k$ candidates for a given query. NDCG@$k$ considers the order of ground-truth objects in the retrieved objects, reflecting the quality of the ranking in the retrieved list. We assign the graded relevance of the result at position $i$ as $rel_{i}\in\{0,1\}$, where $rel_{i}=1$ when the object is relevant to the query, otherwise $rel_{i}=0$. More details can be found in [5]. Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
Baselines of Relevance Model. We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models. We also modify the state-of-the-art textual relevance model DPR [38] into ${}_{D}$ to answer TkQs.
[…]
  • PALM [13]: This method employs deep neural networks for query-object spatio-textual relevance.
  • DrW [5]: This is the newest deep relevance based method for answering TkQs.
  • DPR [38]: This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
[…]
  • MGeo [12]: This is also the newest deep relevance based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper, and therefore we use the evaluation results from the original paper.
Baselines of LIST. We compare LIST with the following retrieve and rerank methods. We select the IR-Tree [2] to accelerate the TkQ search, serving as an index baseline. There exist two categories of ANNS indexes, as discussed in Section II, and our method belongs to the first category. We select the state-of-the-art ANNS indexes in the first category, i.e., IVF [46], LSH [63], and HNSW [52], as index baselines. For the second category of methods, such as product quantization based methods, their standalone efficiency is no better than a brute-force search [^2] https://github.com/facebookresearch/faiss/issues/148. Consequently, following previous work [46], we choose IVFPQ, which integrates PQ with the IVF index from the first category, which is considered as a state-of-the-art solution. Additionally, existing ANNS indexes consider only textual embedding but not the spatial factor. To integrate the spatial factor, we use ${}_{D}$, which is the most effective baseline of the spatio-textual relevance model, to rerank the objects retrieved by an index.
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by the ${}_{D}$. Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above. We evaluate the effect of $k$ in Section V-C.
  • TkQ+${}_{D}$: This method employs TkQ to retrieve top-$k$ objects and then reranks these objects by ${}_{D}$.
  • IVF+${}_{D}$: This method constructs an IVF index over the embeddings produced by ${}_{D}$, and the objects within the selected index cluster are reranked by ${}_{D}$. Notably, this method requires two parameters: the number of clusters $c$, and the number of clusters to route for each query and object $cr$. It does not involve $k$. We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by the ${}_{D}$. Following previous work [63], we set the length of the hash code $nbits$ to 128.
  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by the ${}_{D}$. Following previous work [52], we set the number of links $M$ to 48 and $efConstruction$ to 100.
  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique [46] to retrieve objects. The retrieved objects are reranked by the ${}_{D}$. We set the number of clusters $c$ to be the same as our index. Following the instruction [63], we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4. Here, we do not set $cr$ the same as our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
Due to the page limitation, we demonstrate the effect of key parameters for all these indexes in our full version paper [^3] https://github.com/Heisenberg-Yin/LIST.
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^4] https://github.com/onnx/onnx. In our model, we utilize the bert-base-Chinese pre-trained model [^5] https://huggingface.co/bert-base-chinese from the huggingface Library [66] as encoders, which is the same as previous work [5, 12]. The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets. $\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section V-C. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library [63], while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
TABLE III: Comparison of retrievers across three datasets.
Beijing Shanghai Geo-Glue
Recall NDCG Recall NDCG Recall NDCG
@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
TkQ+${}_{D}$
0.7263 0.6633 0.5056 0.4145 0.7717 0.7348 0.5742 0.4556 0.7357 0.7012 0.5797 0.4878
IVF+${}_{D}$
0.5461
0.4954
0.3771
0.3103
0.4965
0.4717
0.3713
0.3000
0.4720
0.4535
0.3835
0.3274
LSH+${}_{D}$
0.5473
0.5057
0.3973
0.3385
0.5356
0.5097
0.4135
0.3453
0.6686
0.6433
0.5445
0.4661
HNSW+${}_{D}$
0.6267
0.5828
0.4545
0.3765
0.5527
0.5270
0.4182
0.3432
0.5009
0.4807
0.4066
0.3468
IVFPQ+${}_{D}$
0.7064
0.6400
0.4872
0.4002
0.7184
0.6765
0.5257
0.4188
0.6997
0.6724
0.5650
0.4791
LIST 0.7711 0.7170 0.5668 0.4812 0.7721 0.7401 0.6099 0.5139 0.8250 0.7909 0.6747 0.5815
5.63%
7.95%
10.70%
15.32%
0.05%
0.72%
6.21%
12.54%
12.13%
12.79%
16.38%
19.21%
V-B Effectiveness Study (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table II reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW and ${}_{D}$. This can be attributed to the word mismatch issue discussed in Section III-B. Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by [5]. (3) Directly applying the dual-encoder model to our task is less effective. On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance. DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions [12].
Effectiveness of LIST. We investigate the effectiveness of LIST by comparing it with the state-of-the-art baselines. The evaluation results are reported in Table III. We have the following findings: (1) Directly applying ANNS indexes in TkQ task results in severe effectiveness degradation. These inferior results can be attributed to the ignorance of the spatial factor by these ANNS indexes. Compared with other ANNS indexes-based methods, TkQ+${}_{D}$ offers better effectiveness across the three datasets. (2) Overall, LIST significantly outperforms all the baselines, achieving the highest scores on all datasets. Compared with the strongest baseline TkQ+${}_{D}$, LIST achieves up to 12.13% improvements on Recall@10 metric and 19.20% on NDCG@1 metric.
TABLE IV: Query runtime on three datasets (ms).
Beijing
Shanghai
Geo-Glue
TkQ
6.3
6.5
1,178.0
TkQ+${}_{D}$
86.3
84.9
5,412.3
IVF+${}_{D}$
2.9 2.9 4.0
LSH+${}_{D}$
5.2
5.1
31.1
HNSW+${}_{D}$
7.3
6.5
14.6
IVFPQ+${}_{D}$
3.6
3.7
5.2
LIST 3.0 3.0 5.1
V-C Efficiency Study (RQ2)
Runtime of LIST. We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines. During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory. The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set. We present the total runtime of different methods in Table IV. Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST. However, their effectiveness is much worse than LIST as stated in Table III. Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets. Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets. (3) The TkQ+${}_{D}$ exhibits a significantly slower querying speed than TkQ across three datasets. The reason is that IR-Tree exploits a priority queue, using the bottom object to prune lower-scoring objects. As the queue size (hyperparameter $k$) increases, the IR-Tree’s pruning ability declines, aligning with previous empirical results [2].
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Figure 4: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters ($c$) (up and right is better).
Trade-off Study by varying $k$ and $c$. We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings. Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100. For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter top-$k$ from 40,000, 20,000, 1,$0$, 5,000 to 2,000. This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison. Other hyperparameter settings are fixed. We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results. We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency. This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed. (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly. This is due to the word mismatch issue discussed in Section III-B. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+${}_{D}$ exhibits almost constant results. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Figure 5: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$) (up and right is better).
Trade-off Study by varying $k$ and $cr$. We further investigate the impact of the number of clusters $cr$ to route for queries and objects. We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$. For other baselines, we set hyperparameter top-$k$ from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study. Other hyperparameter settings are fixed. Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5. According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings. Increasing $cr$ can enhance effectiveness, but it also leads to higher latency. In practice, users can adjust this parameter according to their needs.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table V, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
TABLE V: Memory usage on three datasets (MB).
Beijing
Shanghai
Geo-Glue
TkQ+${}_{D}$
719
857
12,037
IVF+${}_{D}$
508
505
8,510
LSH+${}_{D}$
513
511
8,638
HNSW+${}_{D}$
548
545
9,427
IVFPQ+${}_{D}$
508
505
8,520
LIST
508
505
8,515
V-D Effect of Dataset Size (RQ3)
Scalability Study. We further explore the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 6. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.
(a) Scalability Study of LIST-R
(b) Scalability Study of LIST
Figure 6: Scalability study on Geo-Glue dataset.
Effect of Training Dataset Size on Effectiveness. We investigate the impact of training dataset size over LIST and LIST-R. Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process. During testing, we use the complete test dataset. This approach avoids the issue of data leakage. We vary the percentage of objects from 30%, 50%, 80% to 100%. We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results. The results are shown in Figure 7(a). LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval. The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
(a) The Impact over Effectiveness
(b) The Impact over Training time
Figure 7: The impact of training dataset size over effectiveness and training time on Beijing dataset.
In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
Effect of Training Dataset Size on Training Time. We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 7(b). Other datasets exhibit similar trends. We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
V-E Ablation Studies (RQ4)
Index Study. To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset. We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects. The results are reported in Table VI, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings. We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table IV), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1. (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed. The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table IV.
TABLE VI: Index ablation study on the Geo-Glue dataset.
Recall NDCG
@20 @10 @5 @1
TkQ+LIST-R
0.7890
0.7585
0.6517
0.5639
IVF+LIST-R
0.5381
0.5192
0.4671
0.3960
LSH+LIST-R
0.7665
0.7381
0.6378
0.5555
HNSW+LIST-R
0.7527
0.7231
0.6219
0.5390
LIST 0.8250 0.7909 0.6747 0.5815
Cluster Quality Study. To illustrate the quality of produced clusters by our index, we conduct a comparison study. We use the proposed relevance model to generate embeddings and then employ our index (denoted as LIST-I) and IVF index to produce clusters separately. We present the cluster results of $IF(C)$ and $P(C)$ in Table VII, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
TABLE VII: Comparison of the quality of clusters.
Beijing Shanghai
IF(C)
$P$(C)
IF(C)
$P$(C)
IVF
1.31
0.6774
1.33
0.6418
LIST-I
1.49
0.8907
1.43
0.8382
Pesudo-Label Study. As discussed in Section IV-C, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index. To investigate the impacts of the pseudo-negative labels, we vary the hyperparameter $\text{neg}_{\text{start}}$ from 40,000, 50,000, 60,000, and 70,000, to 80,000 on the Beijing and Shanghai datasets. We illustrate the metrics $P(C)$ and $\text{IF}(C)$ of produced clusters in Figure 8. Notably, as $\text{neg}_{\text{start}}$ increases, both $IF(C)$ and $P(C)$ tend to increase. An increased $IF(C)$ suggests a more concentrated distribution of objects, while an increased $P(C)$ indicates improved accuracy in the retrieval results. The experiment results indicate that the choice of $\text{neg}_{\text{start}}$ leads to a trade-off between effectiveness and efficiency, which can be set flexibly in real-world applications.
(a) The impact over $IF(C)$
(b) The impact over $P(C)$
[…]
TABLE VIII: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
Recall NDCG
@20 @10 @5 @1
LIST-R + $S_{in}$
0.7526
0.7087
0.5255
0.4271
LIST-R + $\alpha*S_{in}^{\beta}$
0.5308
0.4532
0.3130
0.2411
LIST-R + ADrW
0.7925
0.7414
0.5832
0.4792
LIST-R
0.8156 0.7545 0.5913 0.4989
Weight Learning Study. We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW [5] (denoted as ADrW). Table VIII reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models. Here, LIST-R represents our weight learning mechanism, while LIST-R+ADrW denotes replacing it with the ADrW for training. The results indicate that our weight learning mechanism surpasses the latest ADrW mechanism.
VI CONCLUSIONS AND FUTURE WORK
In this paper, we propose a novel retriever LIST, which comprises a new relevance model and a novel index. We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art baselines. This work opens up a promising research direction of designing novel ANNS indexes for accelerating the search for embedding based spatial keyword queries. One interesting future direction is integrating our index with product quantization techniques to further expedite the search process. Another potential direction is to extend our proposed index to vector databases for dense vectors without spatial information.

### v2

@@
 Abstract
-With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that evaluates both spatial and textual relevance, have found many real-life applications.
-Existing geo-textual indexes for TkQs use traditional retrieval models like BM25 to compute text relevance and usually exploit a simple linear function to compute spatial relevance, but its effectiveness is limited.
-To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues.
-To the best of our knowledge, there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
+With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that considers both spatial and textual relevance, have found many real-life applications.
+To efficiently handle TkQs, many indexes have been developed, but the effectiveness of TkQ is limited.
+To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues and there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
 To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST).
 LIST is featured with two novel components.
-Firstly, we propose a lightweight and effective relevance model that is capable of learning both textual and spatial relevance.
-Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects.
-Two key challenges in building an effective and efficient index are the absence of high-quality labels and unbalanced clustering results.
+Firstly, we propose a lightweight and effective relevance model that is capable of learning both spatial and textual relevance.
+Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index to accelerate the top-k search process of the proposed relevance model, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects.
+There exist two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results.
 We develop a novel pseudo-label generation technique to address the two challenges.
 Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
-I INTRODUCTION
+1 Introduction
 With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated.
-Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps [1]; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location.
+Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location.
 Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent.
 This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements.
-An example query is to retrieve geo-textual objects with texts relevant to the keywords ‘delicious Italian restaurants’ and located near the query’s location, and the result is a list of $k$ geo-textual objects ranked by a relevance model considering both spatial and textual relevance.
-Spatial keyword queries have applications in many scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions.
-To meet diverse user needs, various spatial keyword queries have been introduced [2, 3, 4].
-Among them, the Top-k KNN Spatial Keyword Query (TkQ[2] retrieves top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance.
-According to the experimental evaluation [5], TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query [4], which retrieves the geo-textual objects containing all the query keywords and rank them according to their spatial relevance (or proximity) to the query location.
-Most of the existing studies on spatial keyword queries focus on the improvement in the efficiency of handling spatial keyword queries.
-As such, many indexes [6] and corresponding query processing algorithms have been developed.
-Motivations.
-The existing indexes [2, 7, 8, 9] designed for the TkQ query typically compute textual relevance using traditional retrieval models such as BM25 [10] and TF-IDF [11], which rely on exact word matching to evaluate textual relevance.
-These traditional models often suffer from the word mismatch issue, i.e., synonyms or paraphrases that consist of completely different tokens may convey the same meanings, which limits their effectiveness (as to be shown in Section III-B).
-For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models, although they are relevant to the query.
-As for spatial relevance, existing indexes [2, 7, 8, 9] for the TkQ query typically compute spatial relevance by a linear function of the distance between the query location and the location of a geo-textual object.
-The implicit assumption behind this practice is that the spatial preference for geo-textual objects is a linear function of distance.
-However, this assumption was not examined previously and we find it does not hold on our real-life datasets.
-From our dataset, we find that the spatial preference for geo-textual objects is not a linear function of distance, but exhibits a significant (non-linear) decrease with the increase of distance (detailed in Section III-B).
-This motivates us to design a more effective spatial relevance module.
-A straightforward approach would be to employ an exponential function.
-However, our experiments indicate that this method is even less effective compared to the linear function (shown in Section V-E).
-To improve effectiveness, several deep learning based relevance models [5, 12, 13] propose to to compute textual relevance based on word embeddings.
-For example, DrW [5] employs embeddings generated from the Bert model, identifying top-k relevant terms from geo-textual object description for each query keyword based on the embeddings, and uses an attention mechanism to aggregate scores of each query keyword.
-Although these approaches could achieve better effectiveness, they have very high query latency for computing the relevance score since they rely on embeddings and neural networks (detailed in Section III-B).
+An example query is to search for a ‘delicious Italian restaurant’ that is close to the user’s hotel.
+Spatial keyword queries have applications in many real-world scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions.
+To meet diverse user needs, various spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 .
+Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance.
+Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (formulated in Section 3.1).
+According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient .
+Most of the existing studies congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 on spatial keyword queries focus on improving the efficiency of handling spatial keyword queries.
+As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
+Although many indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited.
+As discussed earlier, TkQ typically uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue, e.g., synonyms that consist of different tokens may convey the same meanings, which limits their effectiveness (detailed in Section 3.2).
+For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models.
+Another limitation arises from using the linear function of distance to calculate spatial relevance.
+This practice is based on an assumption that the user’s spatial preference for geo-textual objects is a linear function of distance.
+However, this assumption was not examined previously and does not hold on our real-life datasets, where the spatial preference for geo-textual objects exhibits a significant (non-linear) decrease with the increase of distance.
+This mismatch with real-world situations results in limited effectiveness (as to be detailed in Section 3.2).
+To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue.
 To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models.
-Existing geo-textual indexes are developed based on traditional retrieval models such as TF-IDF.
-For instance, IR-Tree [2] augments each R-Tree [14] node with a pseudo-inverted file including the words of all objects inside that node.
-These indexes cannot be used to handle embedding based spatial keyword queries, which represent the semantic meaning of query keywords with a high-dimensional embedding.
-Although Approximate Nearest Neighbor Search (ANNS) indexes [15, 16, 17] are designed for embedding retrieval, none of these ANNS indexes take into account the spatial factor, which is essential for spatial keyword queries.
-Directly using these ANNS indexes for spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section V-B).
-It is an open problem to incorporate spatial factors into ANNS indexes to support embedding based spatial keyword queries.
-Objectives.
-To address the aforementioned issues, this work aims to (1) design an embedding based lightweight yet effective spatio-textual relevance model as the ranking function, and (2) design a novel index to speed up high-dimensional embedding based spatial keyword queries.
-To achieve the two objectives, we propose a novel solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries.
-We denote this retriever by LIST, which is featured with two new components.
-First, LIST develops a lightweight and effective relevance model as the ranking function, which comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module.
-Second, LIST introduces a novel machine learning based ANNS index.
-For both embedding based spatial keyword queries and geo-textual objects, our proposed index utilizes the learning-to-cluster technique, which was developed for image clustering in the field of computer vision, to cluster relevant geo-textual objects and queries together and build an index.
-Existing learning-to-cluster techniques  [18, 19] rely on partial pairwise similar/dis-similar relationships of images for training, where the pairwise labels are easy to obtain for images.
-However, such pairwise positive labels for queries and geo-textual objects in our problem are very sparse and high-quality negative labels are absent, which are important for learning-to-cluster.
-Additionally, when the number of clusters exceeds the number of image classes, the existing learning-to-cluster technique will produce a highly skewed cluster distribution [18], which will hurt the efficiency.
+For example, the recently proposed DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings, identifies top-k relevant terms from object description for each query keyword based on the word embeddings, and utilizes an attention mechanism to aggregate the relevance scores of every query keyword as the textual relevance.
+It also uses the linear function of distance as the spatial relevance.
+The interactions between word embeddings bring very high query latency.
+And it relies on interactions between the word embeddings of queries and objects, which cannot be pre-computed until the query-object pairs are given, therefore there is no efficient index designed for it (detailed in Section 3.2).
+To address the aforementioned limitations, this work aims to (1) design a lightweight yet effective relevance model, and (2) design an index to expedite the top-k search process of the proposed relevance model.
+For the first objective, we propose a lightweight yet effective relevance model for embedding based spatial keyword queries, which capture the semantic meaning of query keywords in one high-dimensional embedding.
+The relevance model comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module.
+The dual-encoder module converts query keywords and object descriptions into two separate embeddings and evaluates textual relevance based on the two embeddings.
+To overcome the limitation of linear distance function, as explained earlier, we propose a new learning-based spatial relevance module, which is capable of learning from real-world datasets to better evaluate spatial relevance.
+Additionally, a weight learning module is proposed to adaptively learn a weight to balance the textual and spatial relevance.
+For the second objective, existing geo-textual indexes are developed based on traditional models, and cannot be used to handle embedding based spatial keyword queries.
+Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries.
+Directly using these ANNS indexes for embedding based spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section 5.2).
+An intuitive idea is to adapt these ANNS indexes to incorporate the spatial factor.
+For instance, the IVF index jegouProductQuantizationNearest2011 clusters embeddings using the K-means algorithm and route queries to a small number of close clusters to reduce the search space.
+However, incorporating spatial factors into the K-means algorithm is challenging as it is hard to set the weight to balance spatial and embedding similarities during index construction.
+Manually setting the weights is not only laborious but also ineffective, resulting in inferior effectiveness (as to be shown in Section 5.2).
+It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
+To fill this gap, we develop a novel machine learning based ANNS index.
+To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones.
+Existing learning-to-cluster studies Hsu16_KCL ; Hsu18_L2C demonstrate that high-quality pairwise similar/dis-similar labels are essential for training.
+Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse , and high-quality negative labels are absent.
+Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built on the cluster results.
 To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels.
-These informative negative labels combined with ground-truth positive (relevant) labels (e.g., click-through or human annotation, detailed in Section V-A) enable our index to learn both spatial and textual information, thereby clustering relevant objects and queries together while separating irrelevant objects and queries.
+Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
+Based on the two proposed techniques, namely the lightweight yet effective relevance model and the machine learning based ANNS index, we propose a new solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries (called LIST).
+LIST employs the proposed relevance model to pre-compute embeddings for geo-textual objects and uses the proposed index to cluster the relevant objects together while separating irrelevant ones.
+Given embedding based spatial keyword queries, LIST routes queries to a subset of clusters to reduce the search space.
+Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (detailed in Section 4).
 The contributions of this work are summarized as follows:
+  • New Relevance Model.
+We develop a lightweight yet effective relevance model.
+It is featured with a dual-encoder module to convert query keywords and object descriptions into two separate embeddings and compute textual relevance based on the two embeddings, a novel learning-based spatial relevance module to learn from real-word datasets to better evaluate the spatial relevance, and a new weight learning mechanism to determine the weights between textual and spatial relevance.
   • New Index.
-We develop a novel machine learning based ANNS index that can speed up embedding based spatial keyword queries.
-Different from existing ANNS indexes, it extends the learning-to-cluster technique to cluster relevant objects and queries together.
+We develop a novel machine learning based ANNS index that is tailored for embedding based spatial keyword queries and accelerating the top-k search process of the proposed deep relevance model.
+Different from existing ANNS indexes, it utilizes the learning-to-cluster technique to cluster relevant objects and queries together while separating the irrelevant ones.
 To build an effective and efficient index, we propose a novel pseudo-label generation approach.
-To the best of our knowledge, this is the first index tailored for embedding based spatial keyword queries.
-Additionally, this is the first geo-textual index to utilize a neural network for retrieval without relying on an explicit tree structure and associated heuristic algorithms.
-  • New Relevance Model.
-We develop a lightweight and effective relevance model.
-It is featured with a novel learning-based spatial relevance module to learn spatial relevance adaptively, a new weight learning mechanism to determine the weights between textual and spatial relevance, and a dual-encoder module to compute textual relevance.
+To the best of our knowledge, this is the first index designed for embedding based spatial keyword queries and deep relevance model.
+Additionally, this is the first geo-textual index that employs neural networks for retrieval without relying on an explicit tree structure.
   • Extensive Experiments.
 We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets.
-Experimental results show that LIST outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
-Our implementation is released at https://github.com/Heisenberg-Yin/LIST.
-II RELATED WORK
+Experimental results show that LIST significantly outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
+2 Related Work
 Spatio-Textual Relevance Models.
-Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries [2, 3, 4] are proposed.
-Among them, Top-k KNN Spatial-Keyword Query (TkQ[2] aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial relevance and text relevance.
-The earlier work [2] computes the textual relevance with traditional retrieval models like BM25 [10] and TF-IDF [11] and exploits a simple linear function for spatial relevance.
+Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient are proposed.
+Among them, Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance.
+Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance.
 However, these traditional methods have limited effectiveness.
-To enhance the effectiveness, several deep learning based methods [13, 5, 12] have been developed for query-POI matching, which is essentially spatial keyword query.
-PALM [13] considers geographic information by using location embedding techniques and combines it with semantic representations for query-POI matching.
-DrW [5] computes the deep textual relevance on the term level of query and geo-textual objects and uses the attention mechanism to aggregate scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual relevance and spatial relevance.
-MGeo [12] employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality.
+To enhance the effectiveness, several deep learning based methods DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 have been developed for query-POI matching, which is essentially spatial keyword query.
+PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 considers geographic information by using location embedding techniques and combines it with textual word semantic representations for query-POI matching.
+DrW liu2023effectiveness computes the deep textual relevance on the term level of query keywords and object descriptions.
+It uses the attention mechanism to aggregate the scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual and spatial relevance.
+MGeo DBLP:conf/sigir/DingCXHLZX23 employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality.
 MGeo aligns these two modalities into the same latent space and computes relevance scores based on the produced representations.
-However, these deep learning based methods focus on improving the effectiveness, but ignore the efficiency issue.
+However, these deep learning based methods focus on improving the effectiveness but ignore the efficiency issue.
+Different from previous spatio-textual relevance models that rely on word embeddings and complex models to compute textual relevance, our proposed model employs a dual encoder module to transform query keywords and object descriptions into two embeddings, computes textual relevance based on two embeddings, and we also propose a new learning-based spatial module to learn from real-world datasets to better estimate spatial relevance.
 Spatio-Textual Indexes.
-Although various spatio-textual indexes [3, 4, 7, 2, 8, 9, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] have been designed to efficiently answer spatial keyword queries, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for high-dimensional embedding based retrieval models used by deep learning based methods.
+Various spatio-textual indexes DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:conf/icde/FelipeHR08 ; congEfficientRetrievalTopk2009 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 ; DBLP:conf/icde/ZhangZZL13 ; tao2013fast ; vaid2005spatio ; DBLP:conf/cikm/GobelHNB09 ; DBLP:conf/cikm/ZhouXWGM05 ; DBLP:conf/sigmod/LuLC11 ; DBLP:conf/edbt/ZhangTT13 ; DBLP:conf/icde/ZhangOT10 ; DBLP:conf/icde/ZhangCMTK09 ; DBLP:journals/pacmmod/Sheng0F00C023 ; DBLP:conf/gis/Liu022 have been designed to efficiently answer spatial keyword queries.
+However, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for accelerating the top-k search process of these deep learning based methods.
 […]
-Different from previous spatio-textual indexes developed to accelerate the search for top-k objects with traditional relevance models (which is called sparse retrieval), we aim to design an index to accelerate the search for top-k objects with embedding based relevance models (which is called dense retrieval), which is still an open problem for geo-textual data.
+Additionally, there exists a learned geo-textual index called WISK DBLP:journals/pacmmod/Sheng0F00C023 , which utilizes query workloads to partition geo-textual data to build a tree-based index and employ reinforcement learning techniques to optimize the index.
+However, WISK is designed for spatial keyword range queries, which treats query keywords and query region as Boolean filters, only retrieves objects containing all the query keywords within the given query region, and is not designed to support other queries like the TkQ, which uses keywords to compute textual relevance.
+Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model, which relies on embeddings to compute textual relevance.
+It is still an open problem for geo-textual data.
 Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes.
 Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques.
-The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models [34].
-The first category of models (e.g., ARC-II [35] and MatchPyramid [36]) calculates the word-level similarity between queries and documents for the final relevance scores.
-DrW [5] belongs to this category.
+The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models guo2020deep .
+The first category of models (e.g., ARC-II DBLP:conf/nips/HuLLC14 and MatchPyramid DBLP:conf/aaai/PangLGXWC16 ) calculates the word-level similarity between queries and documents for textual relevance.
+DrW liu2023effectiveness belongs to this category.
 Although this category of methods may have better effectiveness, these methods are usually computationally expensive.
-The second category of models (e.g., DSSM [37]) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations.
+The second category of models (e.g., DSSM huang2013learning ) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations.
 This category of models used to be less effective than the first category of models.
-However, with the emergence of the dual-encoder architecture [38, 39], which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval [40, 41, 42, 43, 44].
-For the second category of methods, to support efficient retrieval, the learned embeddings are usually indexed for Approximate Nearest Neighbor Search (ANNS) [34].
-ANNS indexes are developed to expedite the top-k search on high-dimensional data.
+However, with the emergence of the dual-encoder architecture DBLP:conf/emnlp/KarpukhinOMLWEC20 ; DBLP:conf/icml/GuuLTPC20 , which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 .
+For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep .
+Note that, models in the first category like DrW liu2023effectiveness are not efficient for online computation and retrieval since the complex word interaction function cannot be pre-calculated until we see the input query-object pairs according to guo2020deep .
+ANNS indexes are developed to expedite the top-k search on high-dimensional embeddings.
 These techniques can be broadly categorized into two types.
-The first type focuses on searching a subset of database.
-The representative methods include inverted file index (IVF) based methods [45, 46, 47], hashing-based methods [48, 49, 50, 51], and graph-based methods [52, 17].
-These ANNS indexes rely on heuristic algorithms or centroids to find candidates.
-Specifically, Inverted File Index (IVF) [46] partitions data into clusters using the k-means algorithm and routes queries to clusters based on the query’s distance to the clusters’ centroids.
-Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index [52] construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings.
-Hashing-based methods like Locality-Sensitive Hashing (LSH) index [53] generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates.
-The other type aims to accelerate the search process itself, such as quantization-based methods [46, 54, 55, 56, 54].
+The first type focuses on searching a subset of the database.
+The representative methods include inverted file index (IVF) based methods DBLP:journals/pami/BabenkoL15 ; jegouProductQuantizationNearest2011 ; wang2020deltapq , hashing-based methods luoSurveyDeepHashing2023 ; liu2014sk ; liu2019lsh ; zheng2020pm , and graph-based methods malkovEfficientRobustApproximate2020 ; DBLP:journals/pvldb/WangXY021 .
+These ANNS indexes typically rely on heuristic algorithms to find a subset of candidates.
+Specifically, Inverted File Index (IVF) jegouProductQuantizationNearest2011 partitions data into clusters using the k-means algorithm and routes queries to a subset of close clusters based on the query’s distance to the clusters’ centroids.
+Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index malkovEfficientRobustApproximate2020 construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings.
+Hashing-based methods like Locality-Sensitive Hashing (LSH) index DBLP:journals/corr/WangSSJ14 generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates.
+The other type aims to accelerate the search process itself, such as quantization-based methods jegouProductQuantizationNearest2011 ; DBLP:journals/pami/GeHK014 ; gao2019beyond ; guoAcceleratingLargeScaleInference2020 ; DBLP:journals/pami/GeHK014 .
 Our method belongs to the first category and is orthogonal to the second category.
-Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatio-textual relevance models.
-In contrast, existing ANNS indexes are only for textual relevance.
-(2) Instead of using heuristic algorithms to generate candidates, our proposed index learns from query-object pairs in a data-driven manner and employs neural networks for data partitioning and query routing.
-Notably, our proposed index utilizes a novel learning-to-cluster technique to group geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
-III PROBLEM STATEMENT AND MOTIVATIONS
-III-A Preliminary
-We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual document description $o.doc$.
+Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatial keyword queries.
+In contrast, existing ANNS indexes are designed for embedding retrieval and do not work well when being adapted to embedding based spatial keyword queries (to be shown in Section 5.2).
+(2) Existing ANNS techniques are based on heuristic methods.
+If we want to incorporate spatial factor into the existing ANNS techniques, it is unavoidable to balance the weight between spatial and embedding relevance, which is however challenging.
+To avoid this, our proposed index employs neural networks to learn to cluster relevant geo-textual objects together while separating the irrelevant ones, and routes embedding based spatial keyword queries to relevant clusters through the neural networks to accelerate the top-k search process.
+Notably, our proposed index utilizes a novel learning-to-cluster technique to group embedding based geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
+3 Problem Statement and Motivations
+3.1 Preliminary
+We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual description $o.doc$.
 The location description $o.loc$ is a two-dimensional GPS coordinate composed of latitude and longitude.
-The document description $o.doc$ reflects the textual content.
-The Top-k KNN Spatial-Keyword Query (TkQ[2] is defined as follows.
+The textual description $o.doc$ is a document that describes the object.
+The Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 is defined as follows.
 Top-k KNN Spatial-Keyword Query (TkQ): Given a query $q=\langle loc,doc,k\rangle$, where $q.doc$ denotes the query keywords, $q.loc$ is the query location, and $q.k$ is the number of returned objects, we aim to retrieve $k$ objects with the highest relevance scores $ST(q,o)$:
@@
 […]
-III-B Data Analysis and Motivations
+3.2 Data Analysis and Motivations
 (a) Word Mismatch Analysis
@@
 Word Mismatch.
-As discussed earlier, existing indexes [2, 7, 9] utilize traditional ranking methods, such as BM25 [10] and TF-IDF [11], to compute textual relevance.
-These models often suffer from the word mismatch issue, which reduces their effectiveness.
+As discussed in Section 1, TkQ congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:conf/ssd/RochaGJN11 typically uses traditional information retrieval methods such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 to compute textual relevance.
+These models rely on exact word matching to compute textual relevance and thus suffer from the word mismatch issue, which reduces their effectiveness.
+We further illustrate this issue in the real-world dataset that contains query-object ground-truth pairs.
+The details on the data are given in Section 1.
 As shown in Figure 1(a), in the Beijing dataset, nearly 20% of the ground-truth query-object positive (relevant) pairs have no overlap of words.
-Since these traditional models rely on exact word matching to compute text relevance, relevant objects may receive a low or even zero textual relevance from these models.
-For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero.
+This indicates that objects relevant to the query may receive a low or even zero textual relevance from these models.
+For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero, although the two terms convey the same meaning.
 Spatial Relevance.
-Previous studies [2, 5] typically compute spatial relevance by $1-SDist(q.loc,o.loc)$.
-The implicit assumption for the linear formulation is that the user preference for geo-textual objects is a linear function of distance.
-However, this assumption was not examined previously and we found it does not really hold on our real-life datasets.
+TkQ congEfficientRetrievalTopk2009 ; liu2023effectiveness typically computes spatial relevance by $1-SDist(q.loc,o.loc)$.
+The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance.
+However, this assumption was not examined previously and we found it does not hold on our real-life datasets.
 In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$.
 This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario.
+This motivates us to design a new spatial relevance module to better fit the real-world scenario.
+An intuitive solution would be to employ an exponential function.
+However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.5).
 Efficiency.
-To address this issue, several recent studies [5, 13, 57] have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness.
-However, these methods rely on word embeddings and deep neural networks to compute textual relevance, resulting in high querying latency.
-For instance, DrW [5] needs to identify the top-k relevant words from each object for each query word, and then aggregate these word pairs using an attention mechanism to compute the textual relevance.
-On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in [5].
-This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as rerankers for a small number of retrieved objects.
-Incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges.
-On the one hand, the index techniques developed for spatial keyword queries are designed based on the traditional retrieval models such as BM25 and cannot be used to handle embedding-based spatial keyword queries.
-On the other hand, existing ANNS indexes, which are discussed in Section II, do not consider spatial relevance.
-It is an open problem of incorporating spatial relevance into the ANNS indexes to support spatial keyword queries.
-Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section V-B).
-III-C Problem Statement
-To address these issues, we aim to (1) design a lightweight and effective embedding-based spatio-textual relevance model as the ranking function, and (2) develop an index that can accelerate the top-k search process of embedding based spatial keyword queries.
-IV PROPOSED RETRIEVER (LIST)
+To address the word mismatch issue, several recent studies liu2023effectiveness ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; DBLP:conf/sigir/YuanLLZYZX20 have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness.
+However, these methods rely on word embeddings and complex neural networks to compute textual relevance, resulting in high querying latency.
+For instance, DrW liu2023effectiveness identifies the top-k relevant words from each object’s description for each query word based on their word embeddings, and then aggregates the scores of these word pairs using an attention mechanism to compute the textual relevance.
+On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in liu2023effectiveness .
+This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as re-rankers for a small number of objects returned by a retriever.
+Motivations of LIST.
+As discussed in the last paragraph, incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges.
+However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index.
+Therefore, developing a new lightweight yet effective relevance model and a new index to accelerate the top-k search process for spatial keyword queries based on the proposed deep relevance model is in demand.
+For the first objective, we develop a lightweight yet effective deep relevance model for embedding based spatial keyword queries.
+For the second objective, existing geo-textual indexes DBLP:journals/vldb/ChenCCJ21 developed for spatial keyword queries are designed based on traditional retrieval models such as BM25 and cannot be used to handle embedding based spatial keyword queries.
+Existing ANNS indexes, which are discussed in Section 2, do not consider spatial factors.
+Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section 5.2).
+Simply adapting existing ANNS indexes, such as the IVF index, to incorporate spatial factors will cause even more severe effectiveness degradation, as it is hard to manually set a weight to balance spatial and textual similarities (as to be shown in Section 5.2).
+It is still an open problem to design a new ANNS index to support embedding based spatial keyword queries.
+3.3 Problem Statement
+In this study, we aim to (1) design a lightweight yet effective relevance model, and (2) develop an index to expedite the top-k search process of the proposed relevance model.
+4 Proposed Retriever (LIST)
 Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. LIST consists of a relevance model (shown in yellow) and an index (shown in green).
-IV-A Overview
-As discussed in the introduction, LIST is featured with a novel machine learning based ANNS index and a new relevance model.
-Notably, our index is applicable to any textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
+4.1 Overview
+LIST is featured with a new relevance model and a novel machine learning based ANNS index.
+Notably, our proposed index is also applicable to any other textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
 The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase.
@@
 During the indexing phase, each object is partitioned into one cluster.
-During the query phase, the trained index routes each query to a cluster or a subset of clusters that have the highest probabilities.
-In the cluster, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
-IV-B The Proposed Relevance Model
+During the query phase, the trained index routes each query to either a single cluster or a subset of clusters that have the highest probabilities.
+Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
+4.2 The Proposed Relevance Model
 Textual Relevance Learning.
-Inspired by the dual-encoder model’s great success in document retrieval [38], we employ a dual-encoder module to encode both the query keywords and the document into two separate embeddings and calculate textual relevance using the two embeddings, which is used in spatial keyword query for the first time.
-Compared with previous relevance models [5] which rely on word embeddings and complex nerual networks, this approach is more efficient.
-This module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert [58].
-The encoder takes the textual content of a query or object as input, captures interactions between terms by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional dense vector.
+Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two seperate embeddings, and calculate textual relevance by the inner product between the two embeddings.
+Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
+The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert devlin2018bert .
+The encoder takes the textual content of the query or object as input, captures interactions between words by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding.
 This process is formulated as:
@@
 where $E_{o}(\textperiodcentered;\theta_{o})$ denotes the object encoder parameterized with $\theta_{o}$ and $E_{q}(\textperiodcentered;\theta_{q})$ denotes the query encoder parameterized with $\theta_{q}$.
-Then the text relevance score can be calculated by the inner product between the $q.emb$ and $o.emb$,
+Then the text relevance score is calculated by the inner product between the $q.emb$ and $o.emb$,
 $TRel(q.doc,o.doc)=q.emb\cdot o.emb.$ (3)
 Spatial Relevance Learning.
-To fit the real-world scenario described in Section III-B, it is essential to assign higher spatial relevance to nearby objects and lower spatial relevance to distant objects.
-Therefore, we propose a novel learning-based spatial relevance module to better capture spatial relevance, where $S_{in}=1-SDist(q.loc,o.loc)$ is taken as input.
-The user’s preference for geo-textual objects increases as the distance decreases and exhibits a stepwise decline as the distance increases, i.e., users tend to visit nearby objects and become less tolerant as the distance increases.
-To illustrate this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks.
+As discussed in Section 3.2, it is essential to develop a more effective spatial relevance module.
+Therefore, we propose a new learning-based spatial relevance module to learn to estimate spatial relevance.
+As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern.
+Here, we consider that the user’s geographic preference for geo-textual objects increases as the distance decreases, i.e., users tend to visit nearby objects, and exhibits a stepwise decline as the distance increases.
+To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks.
 If the nearest Starbucks is very close, s/he would go and buy coffee.
@@
 If the nearest Starbucks is very far away, s/he would give up this idea.
-Therefore, a new spatial relevance module is designed, which is a learnable monotonically increasing step function, to align with this real-world characteristic.
-Our learning-based spatial relevance module consists of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$.
-Here, $\mathcal{T}$ is used to ensure that the learned function is a step function and $w_{s}$ is employed to learn and estimate the spatial relevance.
-Specifcially, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance.
-$\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value.
+Therefore, the new spatial relevance module is designed to be a learnable monotonically increasing step function, to align with this real-world characteristic.
+Our proposed spatial relevance module takes $S_{in}=1-SDist(q.loc,o.loc)$ as input, consisting of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$.
+Here, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance, which is used to ensure that the learned function is a step function.
+Specifically, $\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value.
 For example, when $t=100$, $\mathcal{T}$ is $[0.0,0.01,\cdots,0.99,1.0]$.
-The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values.
-Note that weights in $w_{s}$ are learned from the training data.
+The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values, which are learned from the training data and used to estimate the spatial relevance.
 When the input $S_{in}$ exceeds the value of $\mathcal{T}[i]$, then the spatial relevance increases by $act(w_{s}[i])$, where $act$ is an activation function to ensure $act(w_{s}[i])$ remain non-negative.
@@
 Adaptive Weight Learning.
-The latest study [5] has showed the importance of weight learning in improving ranking effectiveness.
-We rely on the dual-encoder architecture to obtain embedding of query $q.emb$ to compute textual relevance.
-In addition, our empirical experiments show that the location embedding technique would introduce additional computational costs.
-Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine an optimal weight based on $q.emb$, formulated as follows:
+The recent study liu2023effectiveness has shown the importance of weight learning in improving ranking effectiveness.
+To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords.
+For instance, in the case of detailed query keywords, such as ‘White House, 1600 Pennsylvania Avenue NW, Washington’, giving greater weight to textual relevance better fits the real-world scenario.
+Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
 $w_{st}=MLP(q.emb),\>w_{st}\in\mathbb{R}^{1\times 2}.$ (6)
@@
 To train our relevance model, we employ the contrastive learning strategy.
-Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section V-A).
-For the negative (irrelevant) geo-textual objects, we choose a subset of hard negative objects for training.
-Specifically, following previous work [5], for each training query, in each training epoch, we randomly pick $b$ hard negative samples from this set.
+Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1).
+As for the negative (irrelevant) geo-textual objects, following previous study liu2023effectiveness , in each training epoch, we randomly pick $b$ hard negative samples from the dataset.
 We optimize the loss function as the negative log-likelihood of the positive object:
 $\begin{split}L_{\text{model}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,b}^{-})\\ &=-log\frac{e^{ST(q_{i},o_{i}^{+})}}{e^{ST(q_{i},o_{i}^{+})}+\sum_{j=1}^{b}e^{ST(q_{i},o_{i,j}^{-})}}.\end{split}$ (8)
-In addition, we also utilize the in-batch negatives strategy [59] to further enhance training efficiency.
-IV-C The Proposed Index
-Figure 3 illustrates our proposed index, which takes geo-location and textual embeddings as input.
-It uses a Multi-Layer Perception (MLP) to produce the $c$-cluster probability distribution and employs training strategies to ensure relevant queries and objects have similar probability distributions.
+In addition, we also utilize the in-batch negatives strategy lin2021inbatch to further enhance training efficiency.
+Figure 3: The illustration of the index.
+4.3 The Proposed Index
+In this section, our objective is to develop a new ANNS index that, for each embedding based spatial keyword query, can return a subset of objects that receive high scores from the proposed relevance model for the input query, thereby accelerating the top-k search process.
+To build such an index, our proposed index adopts a machine learning based method.
+Without manually setting the weight between spatial and textual relevance, it learns to cluster relevant queries and objects together while separating the irrelevant ones.
+Specifically, as illustrated in Figure 3, it takes the geo-location and textual embeddings of objects or queries as input and employs a Multi-Layer Perceptron (MLP) to partition objects and route queries into $c$ clusters, and then searches for the top-k relevant objects within the routed cluster.
+To make the index both effective and efficient, we develop a novel pseudo-negative query-object pair generation method.
+These pseudo labels combined with the positive (relevant) query-object pairs ensure that relevant queries and objects are partitioned and routed into the same cluster while separating the irrelevant ones, thereby reducing the search space while maintaining effectiveness.
 Feature Construction.
@@
 $x=[emb,\hat{lat},\hat{lon}].$ (10)
-Here $x$ is the input feature vector for the cluster classifier.
-Although several region embedding techniques exist [60], we find that this simple approach is sufficiently effective.
-Cluster Classifier.
-Our cluster classifier is a lightweight Multi-Perception Layer (MLP).
-Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
+Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP.
+Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
 $Prob=Softmax(MLP(x)),\>Prob\in\left[0,1\right]^{c\times 1},$ (11)
 where $c$ is a hyperparameter that indicates the desired number of clusters we aim to obtain, and $Prob$ represents the predicted $c$-cluster probability distribution.
-The MLP is shared between queries and objects.
 The $c$-cluster probability distribution of object $o$ is denoted as $Prob_{o}$ and that of query $q$ is represented as $Prob_{q}$.
-The cluster classifier actually serves as a relevance classifier, grouping identified relevant queries and objects into the same cluster.
-For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ forms a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
-Figure 3: The illustration of the index.
+Note that the cluster classifier, i.e., the MLP, is shared between queries and objects.
+This setting allows our index to learn the distribution of both queries and objects, subsequently grouping identified relevant queries and objects into the same cluster (as detailed later).
 Pseudo-Label Generation.
 To train the cluster classifier, both positive and negative pairwise labels are required.
-We utilize the ground-truth query-object relevant label as the positive pairwise label for training.
+We utilize the ground-truth query-object relevant label (e.g., click-through or human annotation, detailed in Section 5.1) as the positive pairwise label for training.
 However, we lack high-quality negative pairwise labels.
@@
 Note that positive query-object pairs are excluded as indicated by the filter condition $s(q,o)=0$.
-This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy [44, 61].
-What sets our approach apart from existing studies is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples.
+This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 .
+What sets our approach apart from existing studies DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples.
 This adjustment strikes the trade-off between the effectiveness and efficiency.
@@
 Thus, the choice of $neg_{start}$ strikes a trade-off between effectiveness and efficiency.
-This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques [18, 19].
-Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function [62] to train the MLP, as described below:
+This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques Hsu16_KCL ; Hsu18_L2C .
+Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function Hsu19_MCL to train the MLP, as described below:
 $\begin{split}L_{\text{Index}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,m}^{-})\\ &=log(\hat{s}(q_{i},o_{i}^{+}))+\sum\limits_{j=1}^{m}log(1-\hat{s}(q_{i},o_{i,j}^{-})),\end{split}$ (14)
@@
 Typically, we randomly select one positive object $o_{i}^{+}$ from the positive object set $\text{pos}_{q}$ and $m$ negative objects from the negative object set $\text{neg}_{q}$ in each training epoch.
-As we aim at grouping the relevant pairs of query and object into the same cluster while the irrelevant pairs are into distinct clusters, the queries and objects are jointly incorporated.
-This is because the MLP is shared between objects and queries.
-In this manner, the complex relationships between queries and objects can be well learned in the process of index construction.
-Figure 3 also demonstrates this process, where relevant pairs share similar output and irrelevant ones have different output.
+As described earlier, the MLP is shared between queries and objects.
+Through this training process, for positive query-object pairs, their $Prob_{o}$ and $Prob_{q}$ will have a similar distribution.
+Thereby, relevant pairs of query and object are more likely to be grouped into the same cluster while the irrelevant pairs are grouped into distinct clusters.
+For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ form a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
 Learning to Partition and Route.
 During the indexing phase, each object $o$ is partitioned to the cluster with the highest probability according to $Prob_{o}$.
-Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which can act as an inverted file for these objects.
-Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which are utilized to calculate the relevance score for incoming queries.
+Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which acts as an inverted file for these objects.
+Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which will be utilized to calculate the relevance score for incoming queries.
 During the query phase, a given query $q$ is directed to the cluster that has the highest probability.
-Subsequently, the relevance scores between $q$ and all objects in the cluster are calculated and the top-k objects are selected as result objects.
+Subsequently, the relevance scores between $q$ and all objects within the cluster are calculated and the top-k objects are selected as result objects.
 An alternative way is routing queries (objects) to $cr$ clusters with the highest probabilities based on $Prob_{q}$ ($Prob_{o}$).
@@
 When an object is deleted, we simply remove it from the corresponding cluster.
-Note that most existing ANNS indexes are static [48, 16, 17].
+Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 .
 When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy.
@@
 Cluster Evaluation.
-After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$.
+After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\\ \cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$.
 We use validation queries to evaluate the quality of the clusters.
@@
 Intuitively, a higher $P(C)$ indicates that the index is more effective.
-In addition, we also consider concerns of efficiency.
-For this purpose, we introduce another metric, the Imbalance Factor (IF) [63], which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$.
+In addition, we also take into account efficiency concerns.
+For this purpose, we introduce another metric, the Imbalance Factor (IF) johnson2019billion , which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$.
 The $IF(C)$ is formulated as:
@@
 In our index, with proper hyperparameters, our training process of the cluster classifier is able to obtain high-quality clusters.
-In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section V-E).
-IV-D Procedures and Analyses of LIST
+In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section 5.5).
+4.4 Procedures and Analyses of LIST
 Input: A geo-textual object dataset $D$, a relevance model $R$, a index $I$, training TkQs set $Q_{\text{train}}$, and incoming TkQs set $Q$
@@
 $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset.
-This is because of the even cluster distribution (Verified in Section V-E), which means approximately $\frac{1}{c}$ of the dataset needs to be visited.
+This is because of the even cluster distribution (Verified in Section 5.5), which means approximately $\frac{1}{c}$ of the dataset needs to be visited.
 Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3.
@@
 $O(nd)$ represents the space required for storing pre-computed $o.emb$, and $O(2n)$ denotes the storage cost for geo-locations.
-V EXPERIMENTS
-TABLE I: Datasets Statistics.
+5 Experiments
+Table 1: Datasets Statistics.
 Beijing
@@
   • RQ2: How well does the efficiency of our proposed retriever LIST compare with state-of-the-art methods?
-  • RQ3: How does LIST perform on datasets and training datasets of varying sizes?
+  • RQ3: How does LIST scale with the size of dataset?
   • RQ4: What are the impacts of our proposed index and different modules in the proposed LIST?
-TABLE II: Comparison of relevance models across three datasets by brute-force search.
+5.1 Experimental Setup
+Datasets.
+To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue.
+The Beijing and Shanghai datasets liu2023effectiveness are provided by Meituan, a Chinese retail services platform.
+Users submit a query through the platform, which consists of a query location and a set of keywords.
+Subsequently, the Points of Interest (POIs) that users clicked on are recorded in the query log, and is considered as a ground-truth.
+Note that the click-through data recorded in the search log may be the only feasible way to get a large scale of ground truth data for spatial keyword queries liu2023effectiveness .
+The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 .
+Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 ; DBLP:journals/csur/WuSZXC23 .
+Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries.
+The Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 is released by DAMO Academic.
+The POIs in the Geo-Glue dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts.
+Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
+The release of query log datasets from the industry is highly restricted.
+As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments.
+To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section 5.4)
+Dataset Split.
+The statistics of datasets are stated in Table 1, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects.
+For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy liu2023effectiveness , where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set.
+From the training set, we randomly choose 10% data as the validation set to tune hyperparameters.
+For the Geo-Glue dataset DBLP:conf/sigir/DingCXHLZX23 , we follow the provided splits for training, validation, and test data.
+Table 2: Comparison of relevance models across three datasets by brute-force search.
 Beijing Shanghai Geo-Glue
@@
 11.70%
-V-A Experimental Setup
-Datasets.
-We utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue.
-The Beijing and Shanghai datasets [5] are provided by Meituan, a Chinese retail services platform.
-Users submit queries consisting of location and keywords.
-The clicked Points of Interest (POIs) recorded in the query log are regarded as the ground truth answers to these queries.Note that the click-through data recorded in the search log is widely used as ground-truth for information retrieval evaluation as explicit feedback such as ratings is challenging to collect and implicit feedback such as click-through data can be tracked automatically and is thus much easier to collect for content providers [64].
-Both datasets are anonymized to protect user privacy and more details can be found here [5].
-The Geo-Glue dataset [65, 12] is released by DAMO Academic.
-The POIs in the dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts.
-Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
-Note that, due to privacy considerations, the release of query log datasets from the industry is restricted.
-As a result, to our knowledge, no other datasets containing ground-truth query results have been released, except for the three datasets used in our experiment.
-To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section V-D)
-Dataset Split.
-The statistics of datasets are stated in Table I, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects.
-For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy [5], where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set.
-From the training set, we randomly choose 10% data as the validation set to tune hyperparameters.
-For the Geo-Glue dataset [12], we follow the provided splits for training, validation, and test data.
 Effectiveness metric.
-Following previous studies [12, 5], we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness.
+Following previous studies DBLP:conf/sigir/DingCXHLZX23 ; liu2023effectiveness , we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness.
 Recall@$k$ evaluates the proportion of positive objects contained in the top-$k$ candidates for a given query.
@@
 We assign the graded relevance of the result at position $i$ as $rel_{i}\in\{0,1\}$, where $rel_{i}=1$ when the object is relevant to the query, otherwise $rel_{i}=0$.
-More details can be found in [5].
+More details can be found in liu2023effectiveness .
 Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
@@
 We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models.
-We also modify the state-of-the-art textual relevance model DPR [38] into ${}_{D}$ to answer TkQs.
+We also modify the state-of-the-art textual relevance model DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 into ${}_{D}$ to answer TkQs.
 […]
-  • PALM [13]: This method employs deep neural networks for query-object spatio-textual relevance.
-  • DrW [5]: This is the newest deep relevance based method for answering TkQs.
-  • DPR [38]: This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
+  • PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 : This method employs deep neural networks for query-object spatio-textual relevance.
+  • DrW liu2023effectiveness : This is the newest deep relevance based method for answering TkQs.
+  • DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 : This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
+  • ${}_{D}$: This method replaces the text relevance computation in DrW liu2023effectiveness with DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 .
+  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep relevance based method for answering TkQs.
+Note that we cannot reproduce the experimental results reported in the paper by running the official code, and we use the evaluation results from the original paper.
 […]
-  • MGeo [12]: This is also the newest deep relevance based method for answering TkQs.
-Note that we cannot reproduce the experimental results reported in the paper, and therefore we use the evaluation results from the original paper.
-Baselines of LIST.
-We compare LIST with the following retrieve and rerank methods.
-We select the IR-Tree [2] to accelerate the TkQ search, serving as an index baseline.
-There exist two categories of ANNS indexes, as discussed in Section II, and our method belongs to the first category.
-We select the state-of-the-art ANNS indexes in the first category, i.e., IVF [46], LSH [63], and HNSW [52], as index baselines.
-For the second category of methods, such as product quantization based methods, their standalone efficiency is no better than a brute-force search [^2] https://github.com/facebookresearch/faiss/issues/148.
-Consequently, following previous work [46], we choose IVFPQ, which integrates PQ with the IVF index from the first category, which is considered as a state-of-the-art solution.
-Additionally, existing ANNS indexes consider only textual embedding but not the spatial factor.
-To integrate the spatial factor, we use ${}_{D}$, which is the most effective baseline of the spatio-textual relevance model, to rerank the objects retrieved by an index.
-An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by the ${}_{D}$.
+An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by ${}_{D}$.
 Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query.
 On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above.
-We evaluate the effect of $k$ in Section V-C.
+We evaluate the effect of $k$ in Section 5.3.
   • TkQ+${}_{D}$: This method employs TkQ to retrieve top-$k$ objects and then reranks these objects by ${}_{D}$.
@@
 We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
-  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by the ${}_{D}$.
-Following previous work [63], we set the length of the hash code $nbits$ to 128.
-  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by the ${}_{D}$.
-Following previous work [52], we set the number of links $M$ to 48 and $efConstruction$ to 100.
-  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique [46] to retrieve objects.
-The retrieved objects are reranked by the ${}_{D}$.
+[…]
+  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by ${}_{D}$.
+Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
+  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by ${}_{D}$.
+Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.
+  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects.
+The retrieved objects are reranked by ${}_{D}$.
 We set the number of clusters $c$ to be the same as our index.
-Following the instruction [63], we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4.
-Here, we do not set $cr$ the same as our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs.
+Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4.
+Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs.
 Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
-Due to the page limitation, we demonstrate the effect of key parameters for all these indexes in our full version paper [^3] https://github.com/Heisenberg-Yin/LIST.
-Implementations.
-The relevance model and the index are trained using Pytorch.
-During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^4] https://github.com/onnx/onnx.
-In our model, we utilize the bert-base-Chinese pre-trained model [^5] https://huggingface.co/bert-base-chinese from the huggingface Library [66] as encoders, which is the same as previous work [5, 12].
-The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets.
-$\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset.
-We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset.
-This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness.
-The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section V-C.
-The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library [63], while others are implemented by ourselves.
-The Faiss library is implemented in C++, providing a fair comparison with our index.
-Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
-TABLE III: Comparison of retrievers across three datasets.
+Table 3: Comparison of retrievers across three datasets.
 Beijing Shanghai Geo-Glue
@@
 0.3274
+[…]
+0.4126
+0.3714
+0.3025
+0.2415
+0.4174
+0.3812
+0.2960
+0.2101
+0.3921
+0.3714
+0.3011
+0.2419
 LSH+${}_{D}$
@@
 19.21%
-V-B Effectiveness Study (RQ1)
+Implementations.
+The relevance model and the index are trained using Pytorch.
+During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx.
+In our model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 .
+The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets.
+$\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset.
+We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset.
+This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness.
+The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section 5.3.
+The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves.
+The Faiss library is implemented in C++, providing a fair comparison with our index.
+Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
+5.2 Effectiveness Study (RQ1)
 Effectiveness of Relevance Model LIST-R.
@@
 All relevance models perform a brute-force search over the entire dataset to identify the top-k objects.
-Table II reports the effectiveness of the evaluated methods.
+Table 2 reports the effectiveness of the evaluated methods.
 DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation.
@@
 On the three datasets, TkQ is outperformed by DrW and ${}_{D}$.
-This can be attributed to the word mismatch issue discussed in Section III-B.
-Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by [5].
+This can be attributed to the word mismatch issue discussed in Section 3.2.
+Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by liu2023effectiveness .
 (3) Directly applying the dual-encoder model to our task is less effective.
 On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance.
-DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions [12].
-Effectiveness of LIST.
-We investigate the effectiveness of LIST by comparing it with the state-of-the-art baselines.
-The evaluation results are reported in Table III.
-We have the following findings: (1) Directly applying ANNS indexes in TkQ task results in severe effectiveness degradation.
-These inferior results can be attributed to the ignorance of the spatial factor by these ANNS indexes.
-Compared with other ANNS indexes-based methods, TkQ+${}_{D}$ offers better effectiveness across the three datasets.
-(2) Overall, LIST significantly outperforms all the baselines, achieving the highest scores on all datasets.
-Compared with the strongest baseline TkQ+${}_{D}$, LIST achieves up to 12.13% improvements on Recall@10 metric and 19.20% on NDCG@1 metric.
-TABLE IV: Query runtime on three datasets (ms).
+DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions DBLP:conf/sigir/DingCXHLZX23 .
+[…]
+Table 4: Query runtime on three datasets (ms).
 Beijing
@@
 LIST 3.0 3.0 5.1
-V-C Efficiency Study (RQ2)
+5.3 Efficiency Study (RQ2)
 Runtime of LIST.
@@
 The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set.
-We present the total runtime of different methods in Table IV.
+We present the total runtime of different methods in Table 4.
 Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset.
 (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST.
-However, their effectiveness is much worse than LIST as stated in Table III.
+However, their effectiveness is much worse than LIST as stated in Table 3.
 Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets.
 Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets.
-(3) The TkQ+${}_{D}$ exhibits a significantly slower querying speed than TkQ across three datasets.
-The reason is that IR-Tree exploits a priority queue, using the bottom object to prune lower-scoring objects.
-As the queue size (hyperparameter $k$) increases, the IR-Tree’s pruning ability declines, aligning with previous empirical results [2].
 (a) Beijing Dataset
@@
 Figure 4: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters ($c$) (up and right is better).
-Trade-off Study by varying $k$ and $c$.
+Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters ($c$).
 We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings.
 Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100.
-For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter top-$k$ from 40,000, 20,000, 1,$0$, 5,000 to 2,000.
+For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter the number of objects retrieved (top-$k$) from 40,000, 20,000, 1,$0$, 5,000 to 2,000.
 This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison.
@@
 (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly.
-This is due to the word mismatch issue discussed in Section III-B.
+This is due to the word mismatch issue discussed in Section 3.2.
 Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases.
@@
 Figure 5: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$) (up and right is better).
-Trade-off Study by varying $k$ and $cr$.
-We further investigate the impact of the number of clusters $cr$ to route for queries and objects.
+Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$).
+We further investigate the impact of the number of clusters $cr$ to route queries and objects.
 We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$.
-For other baselines, we set hyperparameter top-$k$ from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study.
+For other baselines, we set hyperparameter the number of objects retrieved (top-$k$) from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study.
 Other hyperparameter settings are fixed.
@@
 The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance.
-The experiment results are presented in Table V, which demonstrates the remarkable memory efficiency of LIST.
+The experiment results are presented in Table 5, which demonstrates the remarkable memory efficiency of LIST.
 Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory.
@@
 This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
-TABLE V: Memory usage on three datasets (MB).
+Table 5: Memory usage on three datasets (MB).
 Beijing
@@
 8,515
-V-D Effect of Dataset Size (RQ3)
+5.4 Scalibility Study (RQ3)
 Scalability Study.
@@
 We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou.
+Since there is no ground-truth data for the relevance between the newly crawled POIs and the existing queries, we only report the efficiency of our proposed retriever LIST and our relevance model LIST-R.
 For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters.
@@
 Figure 6: Scalability study on Geo-Glue dataset.
-Effect of Training Dataset Size on Effectiveness.
-We investigate the impact of training dataset size over LIST and LIST-R.
-Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process.
-During testing, we use the complete test dataset.
-This approach avoids the issue of data leakage.
-We vary the percentage of objects from 30%, 50%, 80% to 100%.
-We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results.
-The results are shown in Figure 7(a).
-LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval.
-The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
-(a) The Impact over Effectiveness
-(b) The Impact over Training time
-Figure 7: The impact of training dataset size over effectiveness and training time on Beijing dataset.
-In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
-Effect of Training Dataset Size on Training Time.
-We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 7(b).
-Other datasets exhibit similar trends.
-We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
-V-E Ablation Studies (RQ4)
+5.5 Ablation Studies (RQ4)
 Index Study.
@@
 We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects.
-The results are reported in Table VI, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings.
-We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table IV), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1.
+The results are reported in Table 6, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings.
+We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table 4), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1.
 (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed.
-The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table IV.
-TABLE VI: Index ablation study on the Geo-Glue dataset.
+The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table 4.
+Table 6: Index ablation study on the Geo-Glue dataset.
 Recall NDCG
@@
 We use the proposed relevance model to generate embeddings and then employ our index (denoted as LIST-I) and IVF index to produce clusters separately.
-We present the cluster results of $IF(C)$ and $P(C)$ in Table VII, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
-TABLE VII: Comparison of the quality of clusters.
+We present the cluster results of $IF(C)$ and $P(C)$ in Table 7, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
+Table 7: Comparison of the quality of clusters.
 Beijing Shanghai
@@
 Pesudo-Label Study.
-As discussed in Section IV-C, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index.
+As discussed in Section 4.3, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index.
 To investigate the impacts of the pseudo-negative labels, we vary the hyperparameter $\text{neg}_{\text{start}}$ from 40,000, 50,000, 60,000, and 70,000, to 80,000 on the Beijing and Shanghai datasets.
@@
 […]
-[…]
-TABLE VIII: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
+Spatial Learning Study.
+To evaluate the learning-based spatial relevance module, we consider the following baseline: (1) LIST-R+$S_{in}$ that replaces the learning-based spatial relevance module with $S_{in}$ for training, where $S_{in}$ = 1 - $SDist(q.loc,o.loc)$; and (2) LIST-R+$\alpha*S_{in}^{\beta}$ which substitutes the spatial relevance module with a learnable exponential function.
+$\alpha$ and $\beta$ are two learnable parameters and are processed to ensure non-negative.
+Table 8 presents the experimental results obtained by conducting a brute-force search on the Beijing dataset using the trained models.
+Notably, LIST-R outperforms LIST-R+$S_{in}$ in all metrics.
+Interestingly, the first variant outperforms the second variant, which suggests that without careful design, a learnable function may be outperformed by a simple approach.
+Table 8: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
 Recall NDCG
@@
 Weight Learning Study.
-We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW [5] (denoted as ADrW).
-Table VIII reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models.
+We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW liu2023effectiveness (denoted as ADrW).
+Table 8 reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models.
 Here, LIST-R represents our weight learning mechanism, while LIST-R+ADrW denotes replacing it with the ADrW for training.
 The results indicate that our weight learning mechanism surpasses the latest ADrW mechanism.
-VI CONCLUSIONS AND FUTURE WORK
+Effect of Training Dataset Size on Effectiveness.
+We investigate the impact of training dataset size over LIST and LIST-R.
+Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process.
+During testing, we use the complete test dataset.
+This approach avoids the issue of data leakage.
+We vary the percentage of objects from 30%, 50%, 80% to 100%.
+We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results.
+The results are shown in Figure 8(a).
+LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval.
+The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
+(a) The Impact over Effectiveness
+(b) The Impact over Training time
+Figure 8: The impact of training dataset size over effectiveness and training time on the Beijing dataset.
+In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
+Effect of Training Dataset Size on Training Time.
+We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 8(b).
+Other datasets exhibit similar trends.
+We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
+6 Conclusions and Future Work
 In this paper, we propose a novel retriever LIST, which comprises a new relevance model and a novel index.

### v3

@@
 To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues and there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
-To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST).
-LIST is featured with two novel components.
-Firstly, we propose a lightweight and effective relevance model that is capable of learning both spatial and textual relevance.
-Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index to accelerate the top-k search process of the proposed relevance model, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects.
-There exist two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results.
+To tackle these issues, we consider embedding based spatial keyword queries, which capture the semantic meaning of query keywords and object descriptions in two separate embeddings to evaluate textual relevance.
+Although various models can be used to generate these embeddings, no indexes have been specifically designed for such queries.
+To fill this gap, we propose LIST, a novel machine learning based Approximate Nearest Neighbor Search index that Learns to Index the Spatio-Textual data.
+LIST utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects.
+There are two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results.
 We develop a novel pseudo-label generation technique to address the two challenges.
-Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
+Additionally, we introduce a learning based spatial relevance model that can integrates with various text relevance models to form a lightweight yet effective relevance for reranking objects retrieved by LIST.
+Experimental results show that (1) our lightweight embedding based relevance model significantly outperforms state-of-the-art relevance models; (2) LIST outperforms state-of-the-art indexes, providing a better trade-off between effectiveness and efficiency.
 1 Introduction
 With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated.
-Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location.
+Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data that contains both text and geographical location, like photos shared on social platforms that provide both textual descriptions and geographic location.
 Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent.
 This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements.
-An example query is to search for a ‘delicious Italian restaurant’ that is close to the user’s hotel.
-Spatial keyword queries have applications in many real-world scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions.
-To meet diverse user needs, various spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 .
-Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance.
-Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (formulated in Section 3.1).
+An example query is to search for a ‘delicious pizza restaurant’ that is close to the user’s location.
+A returned object could be a nearby restaurant named ‘Pizza Palace’.
+Spatial keyword queries have applications in many real-world scenarios such as geographic search engines DBLP:conf/sigmod/ChenSM06 , location-based services DBLP:conf/cikm/ZhouXWGM05 , and local web advertising tailored to specific regions dong2021continuous .
+To meet diverse user needs, various types of spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 .
+Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves the top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance.
+Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (as to be formulated in Section 3.1).
 According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient .
@@
 As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
-Although many indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited.
-As discussed earlier, TkQ typically uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue, e.g., synonyms that consist of different tokens may convey the same meanings, which limits their effectiveness (detailed in Section 3.2).
-For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models.
-Another limitation arises from using the linear function of distance to calculate spatial relevance.
-This practice is based on an assumption that the user’s spatial preference for geo-textual objects is a linear function of distance.
-However, this assumption was not examined previously and does not hold on our real-life datasets, where the spatial preference for geo-textual objects exhibits a significant (non-linear) decrease with the increase of distance.
-This mismatch with real-world situations results in limited effectiveness (as to be detailed in Section 3.2).
+Motivations.
+Despite various indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited.
+As discussed earlier, TkQ uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue DBLP:conf/emnlp/KarpukhinOMLWEC20 ; renRocketQAv2JointTraining2021 ; DBLP:conf/naacl/QuDLLRZDWW21 , e.g., synonyms that consist of different tokens may convey the same or similar meanings, which limits their effectiveness (as to be detailed in Section 3.2).
+For example, consider searching for an ‘Italian restaurant’ on a location-based app like Foursquare.
+Even if a nearby restaurant named ‘Pasta House’ exists, traditional models will not be able to retrieve this result because they cannot semantically match ‘Italian restaurant’ with ‘Pasta House’, it will receive zero textual relevance score.
 To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue.
-To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models.
-For example, the recently proposed DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings, identifies top-k relevant terms from object description for each query keyword based on the word embeddings, and utilizes an attention mechanism to aggregate the relevance scores of every query keyword as the textual relevance.
-It also uses the linear function of distance as the spatial relevance.
-The interactions between word embeddings bring very high query latency.
-And it relies on interactions between the word embeddings of queries and objects, which cannot be pre-computed until the query-object pairs are given, therefore there is no efficient index designed for it (detailed in Section 3.2).
-To address the aforementioned limitations, this work aims to (1) design a lightweight yet effective relevance model, and (2) design an index to expedite the top-k search process of the proposed relevance model.
-For the first objective, we propose a lightweight yet effective relevance model for embedding based spatial keyword queries, which capture the semantic meaning of query keywords in one high-dimensional embedding.
-The relevance model comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module.
-The dual-encoder module converts query keywords and object descriptions into two separate embeddings and evaluates textual relevance based on the two embeddings.
-To overcome the limitation of linear distance function, as explained earlier, we propose a new learning-based spatial relevance module, which is capable of learning from real-world datasets to better evaluate spatial relevance.
-Additionally, a weight learning module is proposed to adaptively learn a weight to balance the textual and spatial relevance.
-For the second objective, existing geo-textual indexes are developed based on traditional models, and cannot be used to handle embedding based spatial keyword queries.
+For instance, DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings and identifies top-k relevant terms from object description for each query keyword based on the word embeddings.
+It then uses an attention mechanism to aggregate the relevance scores between each keyword and its top-k relevant terms, determining the overall textual relevance.
+In our experiment, DrW takes over 7 seconds to answer a query in the Geo-Glue dataset using brute-force search.
+Moreover, to the best of our knowledge, there are no efficient indexes designed to expedite the top-k search process for these deep relevance models.
+Objective and Challenges.
+To this end, we aim to develop an efficient index for deep learning models to answer spatial keyword queries.
+However, these models, such as DrW, often suffer from high latency and it is difficult to design indexes for them due to their complex word interaction functions, which rely on word embeddings to calculate the word-level similarity between query keywords and object descriptions for evaluating textual relevance.
+According to guo2020deep , these functions not only result in high latency but also cannot be pre-calculated until the query-object text pairs are seen, making it challenging to build an index for them.
+To address this, we consider the embedding based spatial keyword queries, which capture the semantic meanings of query keywords and object descriptions into two separate embeddings and evaluate textual relevance based on the two embeddings.
+Many text models can be used for the embedding based spatial keyword queries, ranging from earlier Word2Vec mikolov2013efficient to the recent advancements in pre-trained language models zhao2024dense .
+The challenge lies in developing an index for embedding based spatial keyword queries.
+Existing geo-textual indexes are developed based on traditional models such as BM25 and TF-IDF, which cannot be used to handle embedding based spatial keyword queries.
 Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries.
@@
 It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
-To fill this gap, we develop a novel machine learning based ANNS index.
+To fill this gap, we develop a new machine learning based ANNS index that is applicable to any relevance models for embedding based spatial keyword queries only if query and objects are represented by two embeddings.
 To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones.
@@
 Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse, and high-quality negative labels are absent.
-Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built on the cluster results.
+Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built based on the cluster results.
 To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels.
 Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
-Based on the two proposed techniques, namely the lightweight yet effective relevance model and the machine learning based ANNS index, we propose a new solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries (called LIST).
-LIST employs the proposed relevance model to pre-compute embeddings for geo-textual objects and uses the proposed index to cluster the relevant objects together while separating irrelevant ones.
-Given embedding based spatial keyword queries, LIST routes queries to a subset of clusters to reduce the search space.
-Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (detailed in Section 4).
+We call the proposed index as LIST, which learns to index spatio-textual data for answering embedding based spatial keyword queries.
+LIST employs a relevance model to pre-compute embeddings for geo-textual objects, and learns to index these embeddings.
+Given an embedding-based spatial keyword query, LIST routes the query to a subset of clusters to reduce the search space.
+Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (as to be detailed in Section 4).
+In addition, we improve the spatial relevance for embedding based spatial keyword queries.
+Existing relevance models such as DrW typically use the linear function of distance, assuming that spatial preference decreases linearly with distance.
+However, this assumption was not examined previously.
+We find it does not hold on our real-life datasets, and that the spatial preference exhibits a significant (non-linear) decrease with the increase of distance (as to be detailed in Section 3.2).
+This motivates us to design a new spatial relevance model to fit this real-world pattern better.
+A straightforward approach would be to employ an exponential function of distance.
+However, our experiments indicate that this simple method is even less effective compared to the linear function (as to be shown in Section 5.6).
+Instead, we design a learning based spatial relevance model, which learns from real-world datasets to better evaluate spatial relevance.
+The spatial model is built on two real-world patterns: (1) spatial relevance increases as distance decreases, and (2) spatial relevance increases in a stepwise manner rather than continuously as distance decreases.
+Therefore, we designed the model as a learnable monotonic step function.
+Additionally, we introduce a weight learning module to adaptively learn a weight for embedding based spatial keyword queries to balance the textual and spatial relevance.
 The contributions of this work are summarized as follows:
-  • New Relevance Model.
-We develop a lightweight yet effective relevance model.
-It is featured with a dual-encoder module to convert query keywords and object descriptions into two separate embeddings and compute textual relevance based on the two embeddings, a novel learning-based spatial relevance module to learn from real-word datasets to better evaluate the spatial relevance, and a new weight learning mechanism to determine the weights between textual and spatial relevance.
   • New Index.
@@
 Additionally, this is the first geo-textual index that employs neural networks for retrieval without relying on an explicit tree structure.
+  • New Spatial Relevance Model.
+We develop a novel learning based spatial relevance module, which is capable of learning from real-world datasets to better evaluate the spatial relevance.
+This model can integrate with various textual relevance models for embedding based spatial keyword queries.
   • Extensive Experiments.
-We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets.
-Experimental results show that LIST significantly outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
+We extensively evaluate the effectiveness and efficiency of our solution LIST on three real-world datasets.
+Experimental results show that (1) our lightweight embedding based relevance model significantly outperforms the state-of-the-art relevance models for effectiveness by an improvement up to 31.60% and 59.92% in terms of NDCG@1 and Recall@10, respectively; (2) LIST outperforms existing state-of-the-art indexes on the three datasets, providing a better trade-off between effectiveness and efficiency.
 2 Related Work
 Spatio-Textual Relevance Models.
-Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient are proposed.
-Among them, Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance.
+Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient have been proposed.
+Among them, Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance.
 Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance.
@@
 However, these deep learning based methods focus on improving the effectiveness but ignore the efficiency issue.
-Different from previous spatio-textual relevance models that rely on word embeddings and complex models to compute textual relevance, our proposed model employs a dual encoder module to transform query keywords and object descriptions into two embeddings, computes textual relevance based on two embeddings, and we also propose a new learning-based spatial module to learn from real-world datasets to better estimate spatial relevance.
+Different from previous deep learning based relevance models that rely on word embeddings and complex models to compute textual relevance, we consider embedding based spatial keyword queries, which transform query keywords and object descriptions into two separate embeddings and computes textual relevance based on the two embeddings.
+We also propose a new learning based spatial module to learn from real-world dataset to better estimate spatial relevance.
 Spatio-Textual Indexes.
@@
 However, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for accelerating the top-k search process of these deep learning based methods.
-[…]
+Several indexes DBLP:journals/www/QianXZZZ18 ; DBLP:journals/geoinformatica/ChenXZZLFZ20 have been introduced to incorporate semantic representations into the TkQ scheme.
+For example, S2R-Tree DBLP:journals/geoinformatica/ChenXZZLFZ20 projects the word embeddings to an $m$-dimensional vector using a pivot-based technique ($m$ as low as 2).
+Consequently, it employs R-trees to index objects based on their geo-locations and $m$-dimensional vectors hierarchically.
+When the dimensionality reaches hundreds, which are common for embeddings, such methods are no better than a linear scan due to the curse of dimensionality indyk1998approximate .
 Additionally, there exists a learned geo-textual index called WISK DBLP:journals/pacmmod/Sheng0F00C023 , which utilizes query workloads to partition geo-textual data to build a tree-based index and employ reinforcement learning techniques to optimize the index.
 However, WISK is designed for spatial keyword range queries, which treats query keywords and query region as Boolean filters, only retrieves objects containing all the query keywords within the given query region, and is not designed to support other queries like the TkQ, which uses keywords to compute textual relevance.
-Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model, which relies on embeddings to compute textual relevance.
+Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model that is developed for embedding based spatial keyword queries.
 It is still an open problem for geo-textual data.
@@
 This category of models used to be less effective than the first category of models.
-However, with the emergence of the dual-encoder architecture DBLP:conf/emnlp/KarpukhinOMLWEC20 ; DBLP:conf/icml/GuuLTPC20 , which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 .
+However, with the emergence of Pre-trained Large Language Models (PLMs), which extract global semantic representation from textual content, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 .
 For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep .
@@
 For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero, although the two terms convey the same meaning.
-Spatial Relevance.
-TkQ congEfficientRetrievalTopk2009 ; liu2023effectiveness typically computes spatial relevance by $1-SDist(q.loc,o.loc)$.
-The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance.
-However, this assumption was not examined previously and we found it does not hold on our real-life datasets.
-In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$.
-This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario.
-This motivates us to design a new spatial relevance module to better fit the real-world scenario.
-An intuitive solution would be to employ an exponential function.
-However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.5).
 Efficiency.
@@
 However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index.
-Therefore, developing a new lightweight yet effective relevance model and a new index to accelerate the top-k search process for spatial keyword queries based on the proposed deep relevance model is in demand.
-For the first objective, we develop a lightweight yet effective deep relevance model for embedding based spatial keyword queries.
-For the second objective, existing geo-textual indexes DBLP:journals/vldb/ChenCCJ21 developed for spatial keyword queries are designed based on traditional retrieval models such as BM25 and cannot be used to handle embedding based spatial keyword queries.
-Existing ANNS indexes, which are discussed in Section 2, do not consider spatial factors.
-Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section 5.2).
-Simply adapting existing ANNS indexes, such as the IVF index, to incorporate spatial factors will cause even more severe effectiveness degradation, as it is hard to manually set a weight to balance spatial and textual similarities (as to be shown in Section 5.2).
-It is still an open problem to design a new ANNS index to support embedding based spatial keyword queries.
-3.3 Problem Statement
-In this study, we aim to (1) design a lightweight yet effective relevance model, and (2) develop an index to expedite the top-k search process of the proposed relevance model.
+Therefore, we consider embeddings based spatial keyword queries, which capture the semantic meaning of query keywords and object descriptions into two separate embeddings.
+Then the challenge lies in how to develop an index for embedding based spatial keyword queries as discussed in Introduction.
+Spatial Relevance.
+Previous studies, such as TkQ congEfficientRetrievalTopk2009 and DrW liu2023effectiveness , typically compute spatial relevance by $1-SDist(q.loc,o.loc)$.
+The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance.
+However, this assumption was not examined previously and we found it does not hold on our real-life datasets.
+In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$.
+This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario.
+This motivates us to design a new spatial relevance module to better fit the real-world scenario.
+An intuitive solution would be to employ an exponential function.
+However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.6).
+Problem Statement.
+We aim to develop an ANNS index for deep learning based relevance models designed for embedding based spatial keyword queries, where query keywords and object descriptions are represented as two separate embeddings.
 4 Proposed Retriever (LIST)
-Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. LIST consists of a relevance model (shown in yellow) and an index (shown in green).
+Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. The relevance model is shown in yellow and the index is shown in green.
 4.1 Overview
-LIST is featured with a new relevance model and a novel machine learning based ANNS index.
-Notably, our proposed index is also applicable to any other textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
-The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase.
-During the training phase, we first train our relevance model, and then train our index.
+LIST is featured with a relevance model and a machine learning based ANNS index.
+Notably, our proposed index is applicable to any other relevance model developed for embedding based spatial keyword queries.
+The workflow of LIST is summarized in Figure 2, which has three phases: training, indexing, and query phase.
+During the training phase, we first train the relevance model, and then train our index.
 During the indexing phase, each object is partitioned into one cluster.
@@
 Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
-4.2 The Proposed Relevance Model
+4.2 The Relevance Model
 Textual Relevance Learning.
-Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two seperate embeddings, and calculate textual relevance by the inner product between the two embeddings.
+Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two separate embeddings, and calculate textual relevance by the inner product between the two embeddings.
 Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
-The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert devlin2018bert .
-The encoder takes the textual content of the query or object as input, captures interactions between words by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding.
+The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, each of which is a BERT model devlin2018bert .
+The encoder takes the textual content of the query or object as input, captures interactions between words by the transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding.
 This process is formulated as:
@@
 As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern.
-Here, we consider that the user’s geographic preference for geo-textual objects increases as the distance decreases, i.e., users tend to visit nearby objects, and exhibits a stepwise decline as the distance increases.
+Here, we consider two features of the preference: (1) it increases as the distance decreases, and (2) it exhibits a stepwise decline as the distance increases.
+The first feature is straightforward because users tend to visit nearby objects.
+Then a straightforward solution is to design a monotonically continuous function to learn from query-object positive pairs and predict users’ spatial preferences.
+However, this approach leads to overfitting.
+For example, if most positive pairs within training dataset are distributed within 1km, the continuous function will learn 1km as a boundary, causing high spatial relevance within 1km but a sharp drop beyond it.
+Actually, 1.1km is not significantly different from 1km for users.
+This overfitting results in poor performance due to lack of generalization (as to be shown in Section 5.6).
+To alleviate this issue, we propose the second feature.
 To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks.
@@
 If the nearest Starbucks is very far away, s/he would give up this idea.
-Therefore, the new spatial relevance module is designed to be a learnable monotonically increasing step function, to align with this real-world characteristic.
+Therefore, this pattern aligns with real-world characteristics and alleviates the overfitting issue as the boundaries are manually based on our knowledge.
+In summary, the new spatial relevance module is designed to be a learnable monotonically step function.
 Our proposed spatial relevance module takes $S_{in}=1-SDist(q.loc,o.loc)$ as input, consisting of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$.
@@
 To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords.
-For instance, in the case of detailed query keywords, such as ‘White House, 1600 Pennsylvania Avenue NW, Washington’, giving greater weight to textual relevance better fits the real-world scenario.
+For instance, in the case of detailed query keywords, such as ‘gas station nearby’, giving lower weight to textual relevance better fits the real-world scenario.
+Since there may be many gas stations nearby, users tend to prefer the closest one.
+Conversely, in another scenario, such as searching with keywords ‘Lincoln Memorial in Washington, D.C.’, due to the uniqueness of the keyword, giving greater weight to textual relevance can deliver better results.
 Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
@@
 Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1).
-As for the negative (irrelevant) geo-textual objects, following previous study liu2023effectiveness , in each training epoch, we randomly pick $b$ hard negative samples from the dataset.
+As for the negative (irrelevant) geo-textual objects, previous studies of passage retrieval DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 have shown that hard negative objects can improve retrieval performance.
+Inspired by them, we introduce this mechanism for spatial keyword query, choosing a subset of hard negative objects for training.
+Specifically, for each training query, we first filter out the positive objects and then use TkQ congEfficientRetrievalTopk2009 to retrieve a set of top-ranked objects, which are considered the hard negative set for that query.
+In each training epoch, we randomly pick $b$ hard negative samples from this set.
 We optimize the loss function as the negative log-likelihood of the positive object:
@@
 In addition, we also utilize the in-batch negatives strategy lin2021inbatch to further enhance training efficiency.
+Complexity Analysis.
+Now we analyze the time complexity of the proposed relevance model.
+Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the embeddings of objects are generated in advance.
+The time complexity of a brute-force search of our relevance model over the entire dataset is $O(n(d+2)+d)$.
+Here, $O(nd)$ represents the time cost of Equation 3, $O(2n)$ corresponds to the time cost of Equation 5, and $O(d)$ accounts for the time cost of Equation 6.
+When compared with other relevance models like DrW liu2023effectiveness , which has a time complexity of $O(nl^{q}(l^{o}d+k^{\prime})+l^{q}(2d^{2}+d))$, our relevance model exhibits superior efficiency in terms of time complexity.
+Here, $l^{q}$ represents the number of words in the query keywords, $l^{o}$ represents the number of words in the object description, and $k^{\prime}$ denotes the number of top-k’ relevant terms to find in the object description for each query keyword.
+These hyperparameters, $l^{q}$, $l^{o}$, and $k^{\prime}$, are typically in the dozens.
+Consequently, DrW is significantly slower than our relevance model.
 Figure 3: The illustration of the index.
@@
 We utilize geo-location and textual embedding to construct a consistent input representation for both objects and queries, which is illustrated in Figure 3.
-The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding, which is converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module.
+The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module, which is typically hunderds, e.g., 768.
+It is L2 normalized before being input into the neural network.
 The geo-location $loc=\langle lat,lon\rangle$ are transformed into the following features $\langle\hat{lat},\hat{lon}\rangle$ as below:
@@
 $x=[emb,\hat{lat},\hat{lon}].$ (10)
-Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP.
+Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP. Since the two feature vectors are normalized independently, they will have a similar impact on the neural network.
 Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
@@
 When an object is deleted, we simply remove it from the corresponding cluster.
+The time cost of the insertion operation is equivalent to the inference time of the neural networks, while the time cost of the deletion operation involves scanning the cluster lists and then deleting the object’s id.
+The time costs of both operations are negligible.
 Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 .
 When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy.
-Our index follows this approach.
-Apparently, the time costs of insertion and deletion operations are negligible.
+This is because most ANNS indexes are based on the principle of organizing similar objects together, thereby routing queries to similar blocks to reduce search space.
+The functions determining which cluster each object belongs to are typically static, such as the distance between objects and the centroids of each cluster in an IVF index.
+When the distribution of embeddings changes significantly, these static functions can no longer accurately capture the new distribution, leading to a drop in performance.
+In our approach, we address this issue by only retraining the index, but not the relevance model.
 Cluster Evaluation.
@@
 Consequently, we calculate the relevance score between $q$ and each object $o$ in that cluster by the relevance model, and the top-$k$ objects with the highest scores are returned to answer query $q$ (lines 17-18).
+In practice, the number of objects assigned to each cluster is relatively small, and thus we evaluate all objects within the routed clusters as it is in existing ANNS indexes.
 Complexity Analysis.
 Now we analyze the time and space complexity of LIST.
-Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the number of layers in the MLP of the index is $l$.
+Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, the number of clusters is $c$, and the number of layers in the MLP of the index is $l$.
 The embeddings $o.emb$ of objects are generated in advance.
@@
 This is because of the even cluster distribution (Verified in Section 5.6), which means approximately $\frac{1}{c}$ of the dataset needs to be visited.
-Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3.
-$O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
 The space complexity of LIST is $O((l-1)d+dc)+n(d+2)$.
@@
 5 Experiments
+In this section, we evaluate the effectiveness and efficiency of our proposed solution for answering TkQs by comparing it with state-of-the-art methods on three real-world datasets.
+We aim to answer the following research questions:
+  • RQ1: Does our relevance model outperform existing relevance models in terms of effectiveness?
+  • RQ2: Does our proposed ANNS index achieve a better effectiveness-efficiency trade-off compared to existing indexes?
+  • RQ3: Can our proposed ANNS index be applied to other relevance models designed for embedding based spatial keyword queries?
+  • RQ4: How does LIST scale with the size of dataset?
+  • RQ4: What are the impacts of our proposed index and different modules in the proposed LIST?
 Table 1: Datasets Statistics.
-Beijing
-Shanghai
-Geo-Glue
-Number of Pois
-122,420
-116,859
-2,849,754
-Number of Queries
-168,998
-127,183
-90,000
-Number of Records
-233,343
-182,634
-90,000
-Number of Train Queries
-136,890
-103,019
-50,000
-Number of Val Queries
-15,209
-11,446
-20,000
-Number of Test Queries
-16,899
-12,718
-20,000
-Number of Train Records
-189,027
-148,017
-50,000
-Number of Val Records
-21,034
-16,492
-20,000
-Number of Test Records
-23,282
-18,125
-20,000
-We evaluate the effectiveness and efficiency of our proposed solution in answering TkQs by comparing with state-of-the-art methods on three real-world datasets.
-We aim to answer the following research questions:
-  • RQ1: Does our proposed retriever LIST and our proposed relevance model outperform previous state-of-the-art methods in terms of effectiveness?
-  • RQ2: How well does the efficiency of our proposed retriever LIST compare with state-of-the-art methods?
-  • RQ3: How does LIST scale with the size of dataset?
-  • RQ4: What are the impacts of our proposed index and different modules in the proposed LIST?
+Beijing Shanghai Geo-Glue
+Number of Pois 122,420 116,859 2,849,754
+Number of Queries 168,998 127,183 90,000
+Number of Records 233,343 182,634 90,000
+Number of Train Queries 136,890 103,019 50,000
+Number of Val Queries 15,209 11,446 20,000
+Number of Test Queries 16,899 12,718 20,000
+Number of Train Records 189,027 148,017 50,000
+Number of Val Records 21,034 16,492 20,000
+Number of Test Records 23,282 18,125 20,000
 5.1 Experimental Setup
@@
 To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue.
-The Beijing and Shanghai datasets liu2023effectiveness are provided by Meituan, a Chinese retail services platform.
+Among them, the Beijing and Shanghai datasets liu2023effectiveness are provided a Chinese retail services platform.
 Users submit a query through the platform, which consists of a query location and a set of keywords.
@@
 The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 .
-Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 ; DBLP:journals/csur/WuSZXC23 .
+Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 .
 Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries.
-The Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 is released by DAMO Academic.
-The POIs in the Geo-Glue dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts.
+In the Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 .
+the POIs are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts.
 Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
@@
 As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments.
-To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section 5.4)
+To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs from OpenStreetMap (as to be shown in Section 5.5)
+Table 2: Hyperparameter settings on three datasets.
+Hyperparameters Beijing Shanghai Geo-Glue
+spatial footstep $t$ 1000 1000 1000
+pseudo-label’s $\text{neg}_{\text{start}}$ 50,000 60,000 180,000
+pseudo-label’s $\text{neg}_{\text{end}}$ 55,000 65,000 181,000
+number of cluster $c$ 20 20 300
 Dataset Split.
@@
 @20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
-TkQ
-0.5740
-0.5283
-0.4111
-0.3302
-0.6746
-0.6380
-0.5044
-0.4009
-0.5423
-0.5023
-0.3847
-0.3051
-PALM
-0.3514
-0.3098
-0.2077
-0.1343
-0.4617
-0.4023
-0.2065
-0.1223
-N/A
-N/A
-N/A
-N/A
-DrW
-0.6968
-0.6316
-0.4814
-0.3791
-0.7689
-0.7159
-0.5394
-0.4114
-N/A
-N/A
-N/A
-N/A
-DPR
-0.5078
-0.4269
-0.2892
-0.2209
-0.5143
-0.4161
-0.2505
-0.1721
-0.7626
-0.7249
-0.6014
-0.5079
-${}_{D}$
-0.7555 0.6845 0.5181 0.4215 0.7986 0.7494 0.5769 0.4568 0.7963 0.7565 0.6211
-0.5202
-MGeo
-N/A
-N/A
-N/A
-N/A
-N/A
-N/A
-N/A
-N/A
-0.7049
-N/A
-N/A
-0.5270
-LIST-R
-0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
-7.95%
-10.22%
-14.12%
-18.36%
-4.69%
-5.73%
-11.71%
-18.14%
-5.73%
-6.18%
-6.26%
-11.70%
+TkQ 0.5740 0.5283 0.4111 0.3302 0.6746 0.6380 0.5044 0.4009 0.5423 0.5023 0.3847 0.3051
+PALM 0.3514 0.3098 0.2077 0.1343 0.4617 0.4023 0.2065 0.1223 N/A N/A N/A N/A
+DrW 0.6968 0.6316 0.4814 0.3791 0.7689 0.7159 0.5394 0.4114 N/A N/A N/A N/A
+MGeo N/A N/A N/A N/A N/A N/A N/A N/A 0.7049 N/A N/A 0.5270
+LIST-R 0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
+(Gain) 17.04% 19.45% 22.82% 31.60% 10.63% 19.48% 19.06% 18.14% 19.06% 59.92% 77.72% 11.70%
 Effectiveness metric.
@@
 Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
-Baselines of Relevance Model.
-We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models.
-We also modify the state-of-the-art textual relevance model DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 into ${}_{D}$ to answer TkQs.
-[…]
+Relevance Model Baselines. We compare our relevance model, denoted as LIST-R, with existing state-of-the-art spatio-textual relevance models.
+  • TkQ congEfficientRetrievalTopk2009 : It uses traditional relevance model BM25 DBLP:journals/ftir/RobertsonZ09 to evaluate text relevance and treats $1-SDist(q.loc,o.loc)$ as spatial relevance.
+The weight parameter $\alpha$ is manually tuned from 0 to 1 with a footstep of 0.1, and the best effectiveness is achieved when $\alpha=0.4$ for all three datasets.
   • PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 : This method employs deep neural networks for query-object spatio-textual relevance.
   • DrW liu2023effectiveness : This is the newest deep relevance based method for answering TkQs.
-  • DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 : This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
-  • ${}_{D}$: This method replaces the text relevance computation in DrW liu2023effectiveness with DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 .
-  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep relevance based method for answering TkQs.
-Note that we cannot reproduce the experimental results reported in the paper by running the official code, and we use the evaluation results from the original paper.
-[…]
-An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by ${}_{D}$.
-Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query.
-On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above.
-We evaluate the effect of $k$ in Section 5.3.
-  • TkQ+${}_{D}$: This method employs TkQ to retrieve top-$k$ objects and then reranks these objects by ${}_{D}$.
-  • IVF+${}_{D}$: This method constructs an IVF index over the embeddings produced by ${}_{D}$, and the objects within the selected index cluster are reranked by ${}_{D}$.
+  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep learning based method for answering TkQs.
+Note that we cannot reproduce the experimental results reported in the paper by running the official code, and thus we use the evaluation results from the original paper.
+Index Baselines.
+We compare our proposed index with existing state-of-the-art indexes.
+Specifically, we select the IR-Tree congEfficientRetrievalTopk2009 to accelerate the TkQ search, serving as an index baseline.
+As discussed in Section 2, existing ANNS indexes fall into two categories and our method belongs to the first category.
+We select the state-of-the-art ANNS indexes in the first category, i.e., IVF jegouProductQuantizationNearest2011 , LSH johnson2019billion , and HNSW malkovEfficientRobustApproximate2020 , as index baselines.
+For the second category of methods, such as product quantization based methods, their standalone efficiency is no better than a brute-force search [^2] https://github.com/facebookresearch/faiss/issues/148.
+Therefore, following previous work jegouProductQuantizationNearest2011 , we choose IVFPQ as an index baseline, which integrates PQ with the IVF index from the first category and is considered as a state-of-the-art solution.
+As discussed in Section 3.2, existing ANNS indexes are limited to considering only textual embeddings, neglecting the spatial factor.
+To mitigate this limitation, we extend the IVF index to include both embedding and spatial factors, which is denoted as IVFS index, as an index baseline.
+Similar to LIST, which uses a retrieve and rerank pipeline, we utilize the aforementioned indexes to retrieve a subset of objects and then rerank the retrieved objects with our relevance model LIST-R for a fair comparison.
+The integration of our proposed index and LIST-R is LIST, and the other methods are detailed below.
+  • TkQ+LIST-R: This method employs TkQ to retrieve top-$k$ objects and reranks these objects by LIST-R.
+  • IVF+LIST-R: This method constructs an IVF index over the embeddings produced by LIST-R, and the objects within the selected index cluster are reranked by LIST-R.
 Notably, this method requires two parameters: the number of clusters $c$, and the number of clusters to route for each query and object $cr$.
@@
 We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
-[…]
-  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by ${}_{D}$.
+  • IVFS+LIST-R: This method follows the IVF+LIST-R pipeline but diverges in the clustering approach.
+Instead of only applying k-means to text embeddings, this approach utilizes k-means on the weighted sum of both geo-location and text embedding factors.
+The weight $\alpha$ is manually tuned from 0 to 1 with a footstep of 0.1, and the best effectiveness is achieved when $\alpha=0.9$ for all these datasets.
+  • LSH+LIST-R: This method constructs an LSH index with the embeddings produced by LIST-R, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by LIST-R.
 Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
-  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by ${}_{D}$.
+  • HNSW+LIST-R: This method constructs an HNSW index over the embeddings produced by LIST-R, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by LIST-R.
 Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.
-  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects.
-The retrieved objects are reranked by ${}_{D}$.
+  • IVFPQ+LIST-R: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects.
+The retrieved objects are reranked by LIST-R.
 We set the number of clusters $c$ to be the same as our index.
-Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4.
+Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 2.
 Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs.
 Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
-Table 3: Comparison of retrievers across three datasets.
-Beijing Shanghai Geo-Glue
-Recall NDCG Recall NDCG Recall NDCG
-@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
-TkQ+${}_{D}$
-0.7263 0.6633 0.5056 0.4145 0.7717 0.7348 0.5742 0.4556 0.7357 0.7012 0.5797 0.4878
-IVF+${}_{D}$
-0.5461
-0.4954
-0.3771
-0.3103
-0.4965
-0.4717
-0.3713
-0.3000
-0.4720
-0.4535
-0.3835
-0.3274
-[…]
-0.4126
-0.3714
-0.3025
-0.2415
-0.4174
-0.3812
-0.2960
-0.2101
-0.3921
-0.3714
-0.3011
-0.2419
-LSH+${}_{D}$
-0.5473
-0.5057
-0.3973
-0.3385
-0.5356
-0.5097
-0.4135
-0.3453
-0.6686
-0.6433
-0.5445
-0.4661
-HNSW+${}_{D}$
-0.6267
-0.5828
-0.4545
-0.3765
-0.5527
-0.5270
-0.4182
-0.3432
-0.5009
-0.4807
-0.4066
-0.3468
-IVFPQ+${}_{D}$
-0.7064
-0.6400
-0.4872
-0.4002
-0.7184
-0.6765
-0.5257
-0.4188
-0.6997
-0.6724
-0.5650
-0.4791
-LIST 0.7711 0.7170 0.5668 0.4812 0.7721 0.7401 0.6099 0.5139 0.8250 0.7909 0.6747 0.5815
-5.63%
-7.95%
-10.70%
-15.32%
-0.05%
-0.72%
-6.21%
-12.54%
-12.13%
-12.79%
-16.38%
-19.21%
+An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by LIST-R.
+To ensure a fair comparison, we set $k$ to 5,000 for the baselines that involve $k$ on the Beijing and Shanghai datasets, which is the average number of objects retrieved by our index per query.
+On the geo-glue dataset, $k$ is set to be 30,000 for the same reason above.
+To investigate the capability of LIST to trade off effectiveness and efficiency, we evaluate the effect of the hyperparameter $cr$ and $k$ in Section 5.3.
 Implementations.
@@
 During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx.
-In our model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 .
-The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets.
-$\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset.
+In our proposed relevance model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 .
+The hyperparameter settings of LIST are detailed in Table 2.
+Specifically, the spatial footstep controller $t$ of $\mathcal{T}$ is set to 1000 for all three datasets.
+The hyperparameters to control the generation of pseudo labels, $\text{neg}_{\text{start}}$ and $\text{neg}_{\text{end}}$, are set to 50,000 and 55,000 for the Beijing dataset, 60,000 and 65,000 for the Shanghai dataset, and 180,000 and 181,000 for the Geo-Glue dataset.
 We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset.
 This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness.
-The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section 5.3.
+The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ settings shown in Section 5.3.
 The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves.
 The Faiss library is implemented in C++, providing a fair comparison with our index.
-Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
-5.2 Effectiveness Study (RQ1)
+Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: V100 32GB.
+(a) Beijing
+(b) Shanghai
+(c) Geo-Glue
+Figure 4: The effectiveness-efficiency trade-off results (upper and right is better).
+5.2 Effectiveness of Proposed Relevance Model (RQ1)
 Effectiveness of Relevance Model LIST-R.
@@
 We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric.
-Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10.
+Specifically, LIST-R achieves up to a 31.60% improvement over the best baseline on NDCG@1 and up to a 59.92% improvement on recall@10.
 (2) Traditional ranking methods are less effective than deep relevance models.
-On the three datasets, TkQ is outperformed by DrW and ${}_{D}$.
+On the three datasets, TkQ is outperformed by DrW.
 This can be attributed to the word mismatch issue discussed in Section 3.2.
-Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by liu2023effectiveness .
-(3) Directly applying the dual-encoder model to our task is less effective.
-On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance.
-DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions DBLP:conf/sigir/DingCXHLZX23 .
-[…]
-Table 4: Query runtime on three datasets (ms).
-Beijing
-Shanghai
-Geo-Glue
-TkQ
-6.3
-6.5
-1,178.0
-TkQ+${}_{D}$
-86.3
-84.9
-5,412.3
-IVF+${}_{D}$
-2.9 2.9 4.0
-LSH+${}_{D}$
-5.2
-5.1
-31.1
-HNSW+${}_{D}$
-7.3
-6.5
-14.6
-IVFPQ+${}_{D}$
-3.6
-3.7
-5.2
-LIST 3.0 3.0 5.1
-5.3 Efficiency Study (RQ2)
-Runtime of LIST.
-We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines.
-During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory.
-The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set.
-We present the total runtime of different methods in Table 4.
-Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset.
-(2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST.
-However, their effectiveness is much worse than LIST as stated in Table 3.
-Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets.
-Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets.
-(a) Beijing Dataset
-(b) Beijing Dataset
-(c) Shanghai Dataset
-(d) Shanghai Dataset
-Figure 4: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters ($c$) (up and right is better).
-Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters ($c$).
-We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings.
-Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100.
-For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter the number of objects retrieved (top-$k$) from 40,000, 20,000, 1,$0$, 5,000 to 2,000.
-This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison.
-Other hyperparameter settings are fixed.
-We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results.
-We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency.
-This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed.
-(2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly.
+Notably, PALM is outperformed by TkQ, which is consistent with the results reported by liu2023effectiveness .
+(a) Beijing
+(b) Shanghai
+Figure 5: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$) (up and right is better).
+5.3 Effectiveness-Efficiency Trade-off of Proposed Index (RQ2)
+Effectiveness-efficiency trade-off of LIST.
+We investigate the effectiveness-efficiency trade-off of LIST by comparing it with the state-of-the-art index baselines, which all use our relevance model.
+The evaluation results are shown in Figure 4, where LIST-R represents the brute-force search using our relevance model.
+We have the following findings: (1) Overall, LIST consistently outperforms all baselines, offering the best trade-off between effectiveness and efficiency across all three datasets.
+Specifically, compared to LIST-R, LIST is an order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset.
+While most ANNS indexes have similar runtime to LIST, their effectiveness falls short.
+(2) Directly applying ANNS indexes to spatial keyword queries results in a significant drop in effectiveness.
+Compared to LIST-R, which uses brute-force search, existing ANNS indexes sacrifice considerable effectiveness for improved efficiency.
+This could be attributed to the fact that these ANNS indexes do not consider the spatial factor.
+(3) When TkQ is used as an index baseline, it is significantly slower.
+The reason is that the IR-Tree is not designed to retrieve a large number of objects.
+As the hyperparameter $k$ increases, the IR-Tree’s pruning ability declines, which is consistent with previous empirical results congEfficientRetrievalTopk2009 .
+(4) Simply modifying the existing ANNS index to incorporate the spatial factor is even less effective.
+IVF+LIST-R consistently outperforms IVFS+LIST-R on the three datasets.
+This confirms that manually assigning a weight to balance the two factors is ineffective.
+Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$).
+We examine LIST’s ability to trade off effectiveness and efficiency.
+Here, we vary the hyperparameter $cr$, which represents the number of clusters to route, from 1, 2, to 3 on the Beijing and Shanghai datasets.
+For baselines that share the same hyperparameter with LIST, i.e., IVF+LIST-R, IVFS+LIST-R, and IVFPQ+LIST-R, we maintain the same increment for $cr$.
+For other methods, we adjust the hyperparameter $k$, which represents the number of objects retrieved by the indexes, to be close to the number of objects retrieved by our proposed index, adjusting it from 5,000, 10,000, to 20,000 on the Beijing and Shanghai datasets.
+For TkQ+LIST-R, due to its slow query processing speed, we adjust $k$ from 100, 500, 1,000, 5,000, and 10,000 to 20,000.
+We present the trade-off results in Figure 5.
+We have the following findings: (1) LIST consistently outperforms other methods, providing a better trade-off between effectiveness and efficiency.
+(2) Compared to other methods, the effectiveness of TkQ+LIST-R increases more slowly as $k$ increases, while the time overhead rises significantly.
 This is due to the word mismatch issue discussed in Section 3.2.
 Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases.
-(3) HNSW+${}_{D}$ exhibits almost constant results.
+(3) HNSW+LIST-R is almost not affected by the number of retrieved objects.
 This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query.
 Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
-(a) Beijing Dataset
-(b) Beijing Dataset
-(c) Shanghai Dataset
-(d) Shanghai Dataset
-Figure 5: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$) (up and right is better).
-Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$).
-We further investigate the impact of the number of clusters $cr$ to route queries and objects.
-We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$.
-For other baselines, we set hyperparameter the number of objects retrieved (top-$k$) from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study.
-Other hyperparameter settings are fixed.
-Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5.
-According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings.
-Increasing $cr$ can enhance effectiveness, but it also leads to higher latency.
-In practice, users can adjust this parameter according to their needs.
 Memory Consumption of LIST.
@@
 The experiment results are presented in Table 4, which demonstrates the remarkable memory efficiency of LIST.
-Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory.
+Compared with LSH+LIST-R, TkQ+LIST-R, and HNSW+LIST-R, LIST requires less memory.
 The reason is that our index stores only a lightweight MLP $c$-cluster classifier.
@@
 Table 4: Memory usage on three datasets (MB).
-Beijing
-Shanghai
-Geo-Glue
-TkQ+${}_{D}$
-719
-857
-12,037
-IVF+${}_{D}$
-508
-505
-8,510
-LSH+${}_{D}$
-513
-511
-8,638
-HNSW+${}_{D}$
-548
-545
-9,427
-IVFPQ+${}_{D}$
-508
-505
-8,520
-LIST
-508
-505
-8,515
-5.4 Scalibility Study (RQ3)
+Beijing Shanghai Geo-Glue
+TkQ+LIST-R 719 857 12,037
+IVF+LIST-R 508 505 8,510
+LSH+LIST-R 513 511 8,638
+HNSW+LIST-R 548 545 9,427
+IVFPQ+LIST-R 508 505 8,520
+LIST 508 505 8,515
+(a) Beijing
+(b) Beijing
+Figure 6: The effectiveness-efficiency trade-off results for different relevance models (upper and right is better).
+5.4 Generalization Study of the Proposed Index (RQ3)
+Generalization Study.
+We explore the applicability of our proposed index to other relevance models designed for embedding based spatial keyword queries.
+Specifically, we introduce two variants of our relevance model: LIST-RRoberta and LIST-RLinear.
+Specifically, in LIST-RRoberta, the BERT model in the dual-encoder module is replaced by another pre-trained language model RoBERTa [^5] https://huggingface.co/clue/roberta_chinese_base wolf2019huggingface .
+In LIST-RLinear, the learned spatial relevance module is replaced with the linear function of distance.
+We train the two baselines on the Beijing dataset and use them to generate embeddings in advance.
+We then use our proposed index to build an index on these embeddings and compare it with existing index baselines.
+The comparison of our proposed index with the baseline indexes is shown in Figure 6.
+The experimental results demonstrate that our proposed index maintains a similar advantage over baseline indexes when applied to other relevance models.
+5.5 Scalibility Study (RQ4)
 Scalability Study.
-We further explore the scalability of LIST and the proposed relevance model LIST-R.
+We evaluate the scalability of LIST and the proposed relevance model LIST-R.
 We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou.
@@
 We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.
-(a) Scalability Study of LIST-R
-(b) Scalability Study of LIST
-Figure 6: Scalability study on Geo-Glue dataset.
+(a) LIST-R
+(b) LIST
+Figure 7: Scalability study on the Geo-Glue dataset.
 5.6 Ablation Studies (RQ4)
-Index Study.
-To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset.
-We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects.
-The results are reported in Table 6, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings.
-We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table 4), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1.
-(2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed.
-The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table 4.
-Table 6: Index ablation study on the Geo-Glue dataset.
-Recall NDCG
-@20 @10 @5 @1
-TkQ+LIST-R
-0.7890
-0.7585
-0.6517
-0.5639
-IVF+LIST-R
-0.5381
-0.5192
-0.4671
-0.3960
-LSH+LIST-R
-0.7665
-0.7381
-0.6378
-0.5555
-HNSW+LIST-R
-0.7527
-0.7231
-0.6219
-0.5390
-LIST 0.8250 0.7909 0.6747 0.5815
 Cluster Quality Study.
@@
 Beijing Shanghai
-IF(C)
-$P$(C)
-IF(C)
-$P$(C)
-IVF
-1.31
-0.6774
-1.33
-0.6418
-LIST-I
-1.49
-0.8907
-1.43
-0.8382
+IF(C) $P$(C) IF(C) $P$(C)
+IVF 1.31 0.6774 1.33 0.6418
+LIST-I 1.49 0.8907 1.43 0.8382
 Pesudo-Label Study.
@@
 @20 @10 @5 @1
-LIST-R + $S_{in}$
-0.7526
-0.7087
-0.5255
-0.4271
-LIST-R + $\alpha*S_{in}^{\beta}$
-0.5308
-0.4532
-0.3130
-0.2411
-LIST-R + ADrW
-0.7925
-0.7414
-0.5832
-0.4792
-LIST-R
-0.8156 0.7545 0.5913 0.4989
+LIST-R + $S_{in}$ 0.7526 0.7087 0.5255 0.4271
+LIST-R + $\alpha*S_{in}^{\beta}$ 0.5308 0.4532 0.3130 0.2411
+LIST-R + ADrW 0.7925 0.7414 0.5832 0.4792
+LIST-R 0.8156 0.7545 0.5913 0.4989
 Weight Learning Study.
@@
 We vary the percentage of objects from 30%, 50%, 80% to 100%.
-We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results.
-The results are shown in Figure 8(a).
-LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval.
+We present the results for NDCG@1 and Recall@10 on the Geo-Glue dataset using default hyperparameters, and similar results are observed for the remaining metrics and datasets.
+The results are shown in Figure 10.
+LIST-R uses brute-force search with the relevance model and LIST utilizes our proposed index for retrieval.
 The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
-(a) The Impact over Effectiveness
-(b) The Impact over Training time
-Figure 8: The impact of training dataset size over effectiveness and training time on the Beijing dataset.
-In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
+Additionally, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
 Effect of Training Dataset Size on Training Time.
-We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 8(b).
+We also measure the training time per epoch of LIST-R and LIST on the Geo-Glue dataset by varying the size of the training data in the same manner stated in the last paragraph, and the results are shown in Figure 9.
 Other datasets exhibit similar trends.
 We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
+Figure 9: The impact of training dataset size on training time for the Geo-Glue dataset.
+Figure 10: The impact of training dataset size on effectiveness on the Geo-Glue dataset.
+Effect of Training Sample Selection on Effectiveness.
+To investigate the impact of training sample selection on effectiveness, we select different 50% subsets from the Geo-Glue training dataset for training and then compare their effectiveness to demonstrate the influence of training sample choice.
+Since the random selection is determined by a random seed, we use different seeds (e.g., 1, 2, 3, 4) to ensure that each selection of training samples is different.
+The experimental results on the Geo-Glue dataset are presented in Figure 11, where we only report the results for NDCG@1, and the other metrics are similar.
+The results show that LIST and LIST-R produce similar results across different training samples, demonstrating that our solution is robust to variations in training sample selection.
+Figure 11: The impact of training sample selection on effectiveness for the Geo-Glue dataset.
+Figure 12: Results of Low-Frequency Descriptors.
+Evaluating Queries Containing Keywords of Low Frequency.
+To investigate the capability of LIST and LIST-R in handling unique or low-frequency keywords, such as ’halal’ or ’vegan’, which are more distinctive than common keywords like ’pizza’ or ’pasta’, we investigate whether the frequency of query keywords affects the performance of LIST and LIST-R on the geo-glue dataset.
+Specifically, we calculate the frequency of each keyword and identify queries containing low-frequency keywords, along with their corresponding ground-truth labels.
+The frequency of query keywords in the geo-glue dataset ranges from 1 to 3,891.
+We select queries containing keywords that appear at most $freq$ times, with $freq$ set to 1, 10, 100, and 1,000, resulting in subsets of test queries consisting of 4,108, 15,892, 19,930, and 19,991 queries, respectively.
+We then evaluate LIST and LIST-R on these subsets to demonstrate their capability in handling queries containing low-frequency keywords.
+The experimental results on the geo-glue dataset are shown in Figure 12, where we only report the results for Recall@10 and NDCG@1, as similar trends are observed on other metrics.
+The results show that LIST and LIST-R maintain performance for queries containing keywords of low-frequency.
 6 Conclusions and Future Work
-In this paper, we propose a novel retriever LIST, which comprises a new relevance model and a novel index.
-We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art baselines.
+In this paper, we propose a novel index LIST for embedding based spatial keyword queries.
+We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art index baselines.
 This work opens up a promising research direction of designing novel ANNS indexes for accelerating the search for embedding based spatial keyword queries.